### v1

# Universal quantum control over non-Hermitian continuous-variable systems

Abstract
Although the control of non-Hermitian quantum systems has a growing interest for their nonunitary feature in the time evolution, the existing discussions are not more than two or three dimensions and heavily influenced by the singularity of the energy spectrum. We here develop a general theory to control an arbitrary number of bosonic modes governed by the time-dependent non-Hermitian Hamiltonian. It takes advantage of the gauge potential in the instantaneous frame rather than the energy spectrum of Hamiltonian. In particular, the dynamics of a general non-Hermitian continuous-variable system is analyzed in the instantaneous frame associated with time-dependent ancillary operators that are superpositions of the laboratory-frame operators and irrelevant to the original Hamiltonian. The gauge potential is determined by the unitary transformation between the time-dependent and stationary ancillary frames. The upper triangularization condition for the Hamiltonian’s coefficient matrix in the stationary ancillary frame enables two of the time-dependent ancillary operators to be nonadiabatic Heisenberg passages of the non-Hermitian system. The probability conservation of the system wavefunction can be restored at the end of these passages without artificial normalization. Our theory is exemplified with the perfect and nonreciprocal state transfers in a cavity magnonic system. The former holds for arbitrary initial states and is irrelevant to the parity-time symmetry of the Hamiltonian and the exceptional point of the spectra; and the latter is consistent with the unidirectional perfect absorbtion. Our work essentially extends the universal quantum control (UQC) theory to the non-Hermitian continuous-variable systems, providing a promising approach for their coherent control.
I Introduction
Conventional control protocols about non-Hermitian systems are spectrum-dependent and typically under the Hamiltonian of parity-time ($\mathcal{PT}$) symmetry Bender and Boettcher (1998); Bender and Hook (2024). Compared to Hermiticity, $\mathcal{PT}$-symmetry of Hamiltonian is a wider criterion for determining whether the system’s eigen-spectra is real or complex. When the Hamiltonian commutes with the parity-time-inversion operator, i.e., $H,\mathcal{PT}]=0$, the system is featured with real eigenenergies. In the broken phase of $\mathcal{PT}$-symmetry, i.e., $H,\mathcal{PT}]\neq 0$, the system can have complex eigenenergies Bender and Boettcher (1998); Bender and Hook (2024). Recently, much effort has been devoted to the presence or absence of exceptional points (EPs) Heiss (2012). In both unbroken and broken phases of $\mathcal{PT}$ symmetry Ashida et al. (2020), EPs might emerge when some eigenvalues and eigenstates of system coalesce and the biorthogonal condition breaks down Brody (2013). The non-Hermitian Hamiltonian arises generally either from the energy or material exchange with the environment Metelmann and Clerk (2015); Harder et al. (2018); Grigoryan et al. (2018); Wang et al. (2019a) or from post-selection over no-quantum-jump trajectories Han et al. (2024). The current work aims to establish a universal theoretical framework for non-Hermitian continuous-variable systems, which is irrelevant to the energy spectrum and is not constrained by the increasing complexity with system size. It is a nontrivial extension of the universal quantum control (UQC) over closed systems Jin and Jing (2025a, b, c, d, e) that are isolated from the external environment.
For the ideal control over closed bosonic systems, the Hermitian Hamiltonian preserves the state norm and the real-valued energy spectrum. However, the quantum systems are inevitably open in practical scenarios, such as photonic systems Guo et al. (2009); Regensburger et al. (2012); Ge and Türeci (2013); Peng et al. (2014); Hodaei et al. (2017); Zhang et al. (2020); Bergman et al. (2021), acoustic systems Zhu et al. (2014); Fleury et al. (2015); Christensen et al. (2016); Liu et al. (2018); Wang et al. (2019b); Tang et al. (2020), and cavity magnonic systems Zhang et al. (2017); Wang et al. (2019a); Zhang et al. (2019); Zhao et al. (2020); Qian et al. (2024); Yang et al. (2024). Their interaction with the environment leads to a breakdown of unitary evolution and a violation of probability conservation of the system wavefunction Ashida et al. (2020). Due to the cooperative coupling to a common reservoir, the non-Hermitian Hamiltonian is used to account for the gain or loss effects associated with individual bosonic modes Zhang et al. (2017, 2025) and the dissipative coupling between the system modes Metelmann and Clerk (2015); Harder et al. (2018); Grigoryan et al. (2018); Wang et al. (2019a). In particular, the loss effect on a cavity mode can be effectively converted into the gain effect by applying external microwave fields to its ports Zhang et al. (2017, 2025).
In the unbroken phase of $\mathcal{PT}$ symmetry and in the absence of EPs, perfect state transfer can be performed in the two- and three-level systems Ibáñez et al. (2011); Torosov et al. (2013, 2014), employing the time-dependent gain or loss rates of populations and the renormalization of the quantum trajectories. By tuning the Rabi frequency and detuning of the driving fields to dynamically encircle an EP, one can create Bell states Khandelwal et al. (2024) and realize the population transfer between two eigenstates Liu et al. (2021) in the nitrogen-vacancy-center system. Within the scattering-matrix framework, a directional quantum amplifier for two photonic modes Metelmann and Clerk (2015) can be achieved when their coherent coupling is precisely matched by their engineered dissipative coupling. Nevertheless, the stringent experimental requirements Ibáñez et al. (2011); Torosov et al. (2013, 2014) and artificial normalization Daley et al. (2009); Uzdin et al. (2012); Ibáñez and Muga (2014); Ashida and Ueda (2018); Dóra and Moca (2020) will become inaccessible when extended to larger-scale systems. The universal quantum control framework Jin and Jing (2025a, b, f, c, g, d, e), however, exploits the gauge potential supported by the differential manifold of rotating frames, on which an energy-spectrum-independent theory can be proposed for the control over a general non-Hermitian bosonic systems.
In this paper, the Heisenberg equation about the non-Hermitian and time-dependent bosonic systems is prescribed by the instantaneous ancillary operators. Transformation from the time-dependent to the time-independent (stationary) ancillary frames introduces a gauge potential, by which the triangularization condition of the Hamiltonian’s coefficient matrix enables the exact solutions of the Heisenberg equation for two of the ancillary operators. In contrast to the conventional treatments of non-Hermitian systems Daley et al. (2009); Uzdin et al. (2012); Ibáñez and Muga (2014); Ashida and Ueda (2018); Dóra and Moca (2020), our theory permits a deterministic evolution toward a desired target state, inherently conserving probability at the end of the passage without artificial normalization. In a cavity magnonic system with gain or loss effects on individual modes and a dissipative coupling between them, it is found that arbitrary initial state in the cavity mode can be perfectly transferred to the magnon mode and perfect absorbtion can occur if the initial excitation is prepared in the magnon mode. In other words, the cavity-magnonic system under our UQC theory can be used as a promising unidirectional perfect absorber Ramezani et al. (2014); Longhi (2015); Jin et al. (2016); Ramezani et al. (2016); Huang et al. (2017); Jin and Song (2018); Xu and Jin (2024) at the operational level.
The rest of this paper is structured as follows. In Sec. II, we introduce a general theory for solving the Heisenberg equation of the bosonic operators under an arbitrary non-Hermitian and time-dependent Hamiltonian. The Hamiltonian’s coefficient matrix can be triangularized in a stationary frame and then it suffices to activate two ancillary operators to become rapid adiabatic passages. In Sec. III.1, our theory is used to analyze a non-Hermitian cavity-magnonic system; and in Sec. III.2, we construct the relevant Heisenberg passages that in general do not conserve the state-norm. Section IV.1 demonstrates that arbitrary state prepared in the cavity mode can be perfectly transferred to the magnon mode irrespective to the parity-time-symmetry of Hamiltonian and the presence or absence of exceptional points. In contrast, Sec. IV.2 shows that a unidirectional perfect absorbtion occurs when the interested state is initially prepared in the magnon mode. The entire work is summarized in Sec. V. Appendix A provides a brief recipe for constructing the ancillary operators as well as the unitary transformation between the time-dependent and stationary ancillary frames. Appendix B details the derivation of the non-Hermitian Hamiltonian based on the Lindblad master equation for the cavity-magnonic system.
II General framework
Consider a general bosonic system consisting of $N$ bosonic modes, associated with the annihilation operators $a_{1}$, $a_{2}$, $\cdots$, $a_{N}$, respectively. The system is controlled by a non-Hermitian Hamiltonian $H(t)$, i.e., $H(t)\neq H^{\dagger}(t)$. Under the assumption of biorthogonal condition Brody (2013), namely that the bra and ket (dual) spaces are equipped with distinct bases, the system dynamics is described by two sets of time-dependent Schrödinger equations as ($\hbar\equiv 1$)
$i\frac{d}{dt}|\psi(t)\rangle$ $=H(t)|\psi(t)\rangle,$ (1a)
$i\frac{d}{dt}|\phi(t)\rangle$ $=H^{\dagger}(t)|\phi(t)\rangle,$ (1b)
where $|\psi(t)\rangle$ and $\langle\phi(t)|$ are the pure-state solutions in the bra and ket spaces, respectively. For the quadratic bosonic systems, the time-dependent Hamiltonian $H(t)$ and its Hermitian-conjugate $H^{\dagger}(t)$ can be expressed as
$H(t)=\vec{a}^{\dagger}H^{a}(t)\vec{a}^{T},\quad H^{\dagger}(t)=\vec{a}^{\dagger}[H^{a}(t)]^{\dagger}\vec{a}^{T},$ (2)
where $\vec{a}\equiv(a_{1},a_{2},\cdots,a_{N})$ and $\vec{a}^{\dagger}=(a_{1}^{\dagger},a_{2}^{\dagger},\cdots,a_{N}^{\dagger})$ are row operator vectors, and $H^{a}(t)$ is an $N\times N$ time-dependent coefficient matrix with Hermitian-conjugate $H^{a}(t)]^{\dagger}$. $H^{a}(t)\neq[H^{a}(t)]^{\dagger}$. The superscript $T$ denotes the matrix transposition, also transforming a row vector to its corresponding column vector. In both discrete- and continuous-variable systems, solving the time-dependent Schrödinger equation with Hermitian or non-Hermitian Hamiltonian remains a challenging task. Here, we extend our UQC theory Jin and Jing (2025a, b, f, c, g, d, e) to the non-Hermitian bosonic systems, providing a fundamental framework for partially solving Eqs. (1a) and (1b).
Without loss of generality, we start with the system dynamics in the ket space, as given by Eq. (1a) with $H(t)$ in Eq. (2). In the spirit of universal quantum control for Hermitian bosonic systems Jin and Jing (2025e), the system dynamics can be described in the ancillary representation associated with a completed set of time-dependent ancillary operators $\mu_{k}(t)$’s, $1\leq k\leq N$. $\{\mu_{k}(t)\}$ are connected with the laboratory-frame bosonic operators $\{a_{k}\}$ by an $N\times N$ unitary transformation matrix $\mathcal{M}^{\dagger}(t)$:
$\vec{\mu}_{t}^{T}=\mathcal{M}^{\dagger}(t)\vec{a}^{T},\quad\vec{\mu}_{t}\equiv[\mu_{1}(t),\mu_{2}(t),\cdots,\mu_{N}(t)].$ (3)
A specific unitary transformation matrix $\mathcal{M}^{\dagger}(t)$ and the relevant formation of ancillary operators are provided in appendix A. Due to Eq. (3) and the unitary feature of $\mathcal{M}^{\dagger}(t)$, one can find that the ancillary operators $\mu_{k}(t)$’s satisfy the canonical communication relation, i.e., $\mu_{j}^{\dagger},\mu_{k}(t)]=-\delta_{jk}$. Using Eq. (3), the time-dependent Hamiltonian (2) can be formulated as
$H(t)=\vec{\mu}_{t}^{\dagger}H^{\mu}(t)\vec{\mu}_{t}^{T},$ (4)
where
$H^{\mu}(t)=\mathcal{M}^{\dagger}(t)H^{a}(t)\mathcal{M}(t)$ (5)
is the Hamiltonian’s coefficient matrix in the representation of time-dependent ancillary bases $\mu_{k}(t)|{\rm vac}\rangle$, where $|{\rm vac}\rangle$ represents the vacuum state.
To proceed, we consider the rotation from the time-dependent ancillary operators to the time-independent or stationary ancillary operators, i.e., $\vec{\mu}_{t}\rightarrow\vec{\mu}_{0}$ with $\vec{\mu}_{0}=[\mu_{1}(0),\mu_{2}(0),\cdots,\mu_{N}(0)]$. Such a rotation can be achieved by the unitary transformation $\mathcal{V}(t)$ as $\mathcal{V}^{\dagger}(t)\mu_{k}(t)\mathcal{V}(t)\rightarrow\mu_{k}(0)$, where $\mathcal{V}(t)$ can be determined by $\mathcal{M}^{\dagger}(t)$ in Eq. (3) (see appendix A for details). Generally, in the rotating frame with respect to $\mathcal{V}(t)$, we have
$H_{\rm rot}(t)$ $=\mathcal{V}^{\dagger}(t)H(t)\mathcal{V}(t)-i\mathcal{V}^{\dagger}(t)\frac{d\mathcal{V}(t)}{dt}$ (6)
$=\vec{\mu}_{0}^{\dagger}\left[H^{\mu}(t)-\mathcal{A}(t)\right]\vec{\mu}_{0}^{T},$
where the dynamical coefficient matrix $H^{\mu}(t)$ is non-Hermitian, i.e., $H^{\mu}(t)\neq[H^{\mu}(t)]^{\dagger}$. The Hermitian and purely geometric matrix $\mathcal{A}$ represents the gauge potential Kolodrubetz et al. (2017) that is associated with the unitary transformation $\mathcal{V}(t)$. The element of the matrix $\mathcal{A}$ in the $k$th row and $m$th column is defined as $\mathcal{A}_{km}=-i[\mu_{k}^{\dagger}(t),d\mu_{m}(t)/dt]$. In parallel, for the bra-space dynamics in Eq. (1b), the Hamiltonian expressed by the stationary ancillary operators is the Hermitian conjugate of $H_{\rm rot}(t)$, i.e., $H^{\dagger}_{\rm rot}(t)=\vec{\mu}_{0}^{\dagger}[(H^{\mu}(t))^{\dagger}-\mathcal{A}(t)]\vec{\mu}_{0}^{T}$.
Consequently, the time-dependent Schrödinger equations in Eq. (1) are transformed as
$i\frac{d}{dt}|\psi(t)\rangle_{\rm rot}$ $=H_{\rm rot}(t)|\psi(t)\rangle_{\rm rot},$ (7a)
$i\frac{d}{dt}|\phi(t)\rangle_{\rm rot}$ $=H^{\dagger}_{\rm rot}(t)|\phi(t)\rangle_{\rm rot},$ (7b)
with the rotated pure states
$|\psi(t)\rangle_{\rm rot}=\mathcal{V}^{\dagger}(t)|\psi(t)\rangle,\quad|\phi(t)\rangle_{\rm rot}=\mathcal{V}^{\dagger}(t)|\phi(t)\rangle.$ (8)
The evolution operators for $|\psi(t)\rangle_{\rm rot}$ and $|\phi(t)\rangle_{\rm rot}$ are
$U_{\rm rot}(t)=\hat{T}e^{-i\int_{0}^{t}H_{\rm rot}(s)ds},\quad V_{\rm rot}(t)=\hat{T}e^{-i\int_{0}^{t}H_{\rm rot}^{\dagger}(s)ds},$ (9)
respectively, where $\hat{T}$ is the time-ordering operator.
Still it remains difficult to determine the system dynamics by solving the evolution operators in Eq. (9), owing to the noncommunity of $H_{\rm rot}(t)$ at distinct moments. And it is not appropriate to apply the commutation condition about the Hamiltonian’s coefficient matrix and the projection operator in the representation of time-independent ancillary modes established for the Hermitian bosonic network Jin and Jing (2025e) to Eq. (6) due to the fact that the non-Hermitian coefficient matrix $H^{\mu}(t)-\mathcal{A}(t)$ is generally non-diagonalizable, as $H^{\mu}_{km}(t)\neq H^{\mu}_{mk}(t)$ for $k\neq m$.
Main result.— we prove that the upper triangularization of the coefficient matrix of $H_{\rm rot}(t)$ is a sufficient condition for activating two useful Heisenberg-picture passages, which is only relevant to the differential manifold rather than the spectral characteristics. Specifically, the upper triangularization condition for the coefficient matrix $H^{\mu}(t)-\mathcal{A}(t)$ is given by Jin and Jing (2025g); Ju et al. (2024)
$\left[H^{\mu}(t)-\mathcal{A}(t)\right]\Pi^{k}-\Pi^{k}\left[H^{\mu}(t)-\mathcal{A}(t)\right]^{T}=0,$ (10)
where $T$ means matrix transposition and $\Pi^{k}$ with $1\leq k\leq N$ is the projection operator or an $N\times N$ matrix defined by $\Pi^{k}_{jm}=\delta_{jk}\delta_{mk}$.
Proof.— The upper triangularization condition given by Eq. (10) yields the upper triangularized Hamiltonian:
$H_{\rm rot}(t)=\sum_{k=1}^{N}\sum_{m\geq k}^{N}\left[H^{\mu}_{km}(t)-\mathcal{A}_{km}(t)\right]\mu_{k}^{\dagger}(0)\mu_{m}(0).$ (11)
The dynamics of an arbitrary operator $\tilde{\mathcal{O}}_{S}$ can be obtained by $\mathcal{O}_{H}(t)=V_{\rm rot}^{\dagger}(t)\tilde{\mathcal{O}}_{S}U_{\rm rot}(t)$ according to the non-Hermitian Heisenberg equation Miao and Xu (2016) with the evolution operators given by Eq. (9). Specifically under the Hamiltonian (11), the dynamics of the ancillary operators $\mu_{1}^{\dagger}(0)$ and $\mu_{N}(0)$ are found to be decoupled from the others and therefore can be derived as $v_{1}^{\dagger}(t)\equiv V_{\rm rot}^{\dagger}(t)\mu_{1}^{\dagger}(0)U_{\rm rot}(t)$ and $v_{N}(t)\equiv V_{\rm rot}^{\dagger}(t)\mu_{N}(0)U_{\rm rot}(t)$. In particular, we have
$\frac{dv_{1}^{\dagger}(t)}{dt}$ $=iV_{\rm rot}^{\dagger}(t)\left[H_{\rm rot}(t),\mu_{1}^{\dagger}(0)\right]U_{\rm rot}(t)$ (12)
$=i\left[H^{\mu}_{11}(t)-\mathcal{A}_{11}(t)\right]V_{\rm rot}^{\dagger}(t)\mu_{1}^{\dagger}(0)U_{\rm rot}(t)$
$=i\left[H^{\mu}_{11}(t)-\mathcal{A}_{11}(t)\right]v_{1}^{\dagger}(t),$
and
$\frac{dv_{N}(t)}{dt}$ $=iV_{\rm rot}^{\dagger}(t)\left[H_{\rm rot}(t),\mu_{N}(0)\right]U_{\rm rot}(t)$ (13)
$=-i\left[H^{\mu}_{NN}(t)-\mathcal{A}_{NN}(t)\right]V_{\rm rot}^{\dagger}(t)\mu_{N}(0)U_{\rm rot}(t)$
$=-i\left[H^{\mu}_{NN}(t)-\mathcal{A}_{NN}(t)\right]v_{N}(t).$
The solutions to the differential equations (12) and (13) are given by
$v_{1}^{\dagger}(t)=e^{if_{1}(t)}\mu_{1}^{\dagger}(0),\quad v_{N}(t)=e^{-if_{N}(t)}\mu_{N}(0),$ (14)
respectively, where the global phases are defined as
$f_{k}(t)=\int_{0}^{t}ds\left[H^{\mu}_{kk}(s)-\mathcal{A}_{kk}(s)\right],\quad k=1,N.$ (15)
The dynamics of the operators in the picture governed by $H_{\rm rot}(t)$ is related to that in the original picture controlled by $H(t)$ through Eq. (6). According to Eqs. (8) and (14), the dynamics of the ancillary operators $\mu_{1}^{\dagger}(0)$ and $\mu_{N}(0)$ in the original picture can be obtained as
$\mu_{1}^{\dagger}(0)$ $\rightarrow\mathcal{V}(t)v_{1}^{\dagger}(t)\mathcal{V}^{\dagger}(t)$ (16)
$=e^{if_{1}(t)}\mathcal{V}(t)\mu_{1}^{\dagger}(0)\mathcal{V}^{\dagger}(t)=e^{if_{1}(t)}\mu_{1}^{\dagger}(t),$
and
$\mu_{N}(0)$ $\rightarrow\mathcal{V}(t)v_{N}(t)\mathcal{V}^{\dagger}(t)$ (17)
$=e^{-if_{N}(t)}\mathcal{V}(t)\mu_{N}(0)\mathcal{V}^{\dagger}(t)=e^{-if_{N}(t)}\mu_{N}(t).$
Equations (16) and (17) indicate that the ancillary operators $\mu_{1}^{\dagger}(t)$ and $\mu_{N}(t)$ are activated as useful passages toward the desired target state. For the systems initially prepared in the states $F[\mu_{1}^{\dagger}(0)]|{\rm vac}\rangle$ and $F[\mu_{N}(0)]|{\rm vac}\rangle$, they will evolve to the states $F[e^{if_{1}(t)}\mu_{1}^{\dagger}(t)]|{\rm vac}\rangle$ and $F[e^{-if_{N}(t)}\mu_{N}(t)]|{\rm vac}\rangle$ at the time $t$ with the accumulated global phases $f_{1}(t)$ and $f_{N}(t)$, respectively, where $F[\cdot]$ represents an arbitrary function of operators.
In parallel, for the systems governed by the Hermitian-conjugate Hamiltonian $H_{\rm rot}^{\dagger}(t)$ in the dual space, the upper triangularization condition (10) can activate $\mu_{1}(t)$ and $\mu_{N}^{\dagger}(t)$ as nonadiabatic passages in the Heisenberg picture. The Hermitian conjugate of Eq. (11) takes the down triangularization form of
$H^{\dagger}_{\rm rot}(t)=\sum_{k=1}^{N}\sum_{m\geq k}^{N}\left[(H^{\mu}_{km}(t))^{*}-\mathcal{A}_{km}(t)\right]\mu_{m}^{\dagger}(0)\mu_{k}(0).$ (18)
Similar to the ket-space case, the dynamics of the ancillary operators $\mu_{1}(0)$ and $\mu_{N}^{\dagger}(0)$ can be obtained as
$\mu_{1}(0)\rightarrow e^{if_{1}^{*}(t)}\mu_{1}(t),\quad\mu_{N}^{\dagger}(0)\rightarrow e^{-if_{N}^{*}(t)}\mu_{N}^{\dagger}(t)$ (19)
by using the non-Hermitian Heisenberg equation and the Hamiltonian (18), where the global phase $f_{k}^{*}(t)$, $k=1,N$, is the complex conjugate of $f_{k}(t)$ in Eq. (15).
III Non-Hermitian cavity-magnonic system
III.1 Model and Hamiltonian
Figure 1: Sketch of an open cavity magnonic system under control. More than the coherent exchange coupling with the strength $J$ and the phase $\varphi$ between the cavity mode $a$ and the magnon mode $b$, the cooperative coupling of the two-mode system to the environment gives rise to the gain or loss rate $\gamma_{a}$ of the cavity mode, the loss rate $\gamma_{b}$ of the magnon mode, and the dissipation coupling $i\Gamma e^{i\Theta}$ between them. $\Theta$ can be 0 or $\pi$ in experiments Wang et al. (2019a).
In this section, we use our UQC theory for the general non-Hermitian bosonic systems in Sec. II to analyze an open two-mode bosonic system that consists of a single-mode cavity and a yttrium iron garnet sphere in the Kittel mode Harder et al. (2018); Grigoryan et al. (2018); Bhoi et al. (2019); Yang et al. (2019); Yao et al. (2019); Yu et al. (2019); Boventer et al. (2020) as shown in Fig. 1. The cavity and magnon modes are represented by the annihilation operators $a$ and $b$, respectively. They are mutually interacted by the time-dependent exchange coherent coupling $J(t)$ with the phase $\varphi$. The cooperative coupling between the two-mode system and the environment Metelmann and Clerk (2015); Wang et al. (2019a) leads to the gain or loss rates $\gamma_{a}$ and $\gamma_{b}$ of the cavity mode $a$ and the magnon mode $b$, respectively, as well as a dissipative coupling $i\Gamma e^{i\Theta}$ between them. Then the whole system can be described by a non-Hermitian Hamiltonian Metelmann and Clerk (2015); Wang et al. (2019a):
$H(t)=(\omega_{a}-i\gamma_{a}e^{i\varphi_{a}})a^{\dagger}a+[\omega_{b}(t)-i\gamma_{b}]b^{\dagger}b$ (20)
$+$ $\left[J(t)e^{i\varphi}+i\Gamma e^{i\Theta}\right]a^{\dagger}b+\left[J(t)e^{-i\varphi}+i\Gamma e^{i\Theta}\right]b^{\dagger}a.$
Here $\omega_{a}$ is the static frequency of the cavity mode. $\omega_{b}(t)$ is the time-dependent frequency of the magnon modes, which can be flexibly tuned by the external bias magnetic field $B(t)$ Zhang et al. (2014); Zare Rameshti et al. (2022); Jin and Jing (2024). $\gamma_{b}>0$ is the fixed loss rate of magnon mode. $\gamma_{a}>0$ represents the gain or loss effect on the cavity mode when $\varphi_{a}=\pi$ or $\varphi_{a}=0$, respectively Zhang et al. (2017, 2025). The relative phase $\Theta=0$ or $\Theta=\pi$ between the dissipative coupling and the coherent coupling is determined by the propagation direction of the traveling waves Wang et al. (2019a). For simplicity, we set $\Theta=0$ in the following. The details about the derivation of the non-Hermitian Hamiltonian (20) with $\varphi_{a}=0$ through the Lindblad master equation can be found in appendix B, despite both $\varphi_{a}$ and $\Gamma$ can have many choices Wang et al. (2019a); Zhang et al. (2017, 2025).
In the rotating frame with respect to $H_{0}(t)=\omega_{0}(t)(a^{\dagger}a+b^{\dagger}b)$, the full Hamiltonian (20) is transformed as
$H(t)$ $=\left[\frac{\Delta(t)}{2}-i\gamma_{a}e^{i\varphi_{a}}\right]a^{\dagger}a-\left[\frac{\Delta(t)}{2}+i\gamma_{b}\right]b^{\dagger}b$ (21)
$+\left[J(t)e^{i\varphi}+i\Gamma\right]a^{\dagger}b+\left[J(t)e^{-i\varphi}+i\Gamma\right]b^{\dagger}a,$
where the detuning $\Delta(t)$ follows $\Delta(t)/2=\omega_{a}-\omega_{0}(t)=-\omega_{b}(t)+\omega_{0}(t)$. One can find that the system Hamiltonian (21) satisfies $\mathcal{PT}$-symmetry Ashida et al. (2020); Bender and Hook (2024) under the setting of $\varphi_{a}=\pi$, $\gamma_{a}=\gamma_{b}$, $\Delta(t)=0$, and the vanishing dissipative coupling $\Gamma=0$. In fact, the Hamiltonian eigenenergies
$E_{\pm}=-i\frac{\gamma_{a}e^{i\varphi_{a}}+\gamma_{b}}{2}\pm\frac{1}{2}\Big\{\Delta^{2}-\left(\gamma_{a}e^{i\varphi_{a}}-\gamma_{b}\right)^{2}$ (22)
$+4(J^{2}-\Gamma^{2})+i\left[8J\Gamma\cos\varphi-2\Delta\left(\gamma_{a}e^{i\varphi_{a}}-\gamma_{b}\right)\right]\Big\}^{1/2}.$
become real-valued, i.e., $E_{\pm}=\pm\sqrt{J^{2}-\gamma_{a}^{2}}$, when $H(t),\mathcal{PT}]=0$ and $J>\gamma_{a}$. Otherwise, they remain complex and the Hamiltonian is in the $\mathcal{PT}$-symmetry-broken phase. Whether holding $\mathcal{PT}$-symmetry or not, the system can approach EP under the conditions of $\Delta(t)=0$, $\varphi=\pi/2$, $\gamma_{a}=\gamma_{b}$, and $J=\pm\Gamma$ (when $\varphi_{a}=0$) or $J=\pm\sqrt{\Gamma^{2}+\gamma_{a}^{2}}$ (when $\varphi_{a}=\pi$), where the eigenenergies coalesce as $E_{+}=E_{-}$.
Conventionally, the control protocols proposed for non-Hermitian systems are based on their spectrum characteristics Ashida et al. (2020); Bender and Hook (2024), focusing on $\mathcal{PT}$-symmetry and EPs. In contrast, our UQC theory in Sec. II is essentially proposed in the instantaneous frames, which are irrespective to the spectral properties. In the following section, we will demonstrate that an arbitrary state can be perfectly transferred from the cavity mode to magnon mode, regardless of whether the system in the unbroken or broken phases of $\mathcal{PT}$-symmetry and whether EPs are present or not.
III.2 Universal Heisenberg passages of two operators
With Eq. (21), we construct the non-unitary universal passages for the non-Hermitian cavity-magnonic system, where the final state can be automatically normalized. According to Eqs. (42) and (43), the time-dependent ancillary operators of an arbitrary two-mode system can be written as
$\mu_{1}(t),\mu_{2}(t)]^{T}=\mathcal{M}^{\dagger}(t)(a,b)^{T}$ (23)
with a $2\times 2$ unitary transformation matrix
$\mathcal{M}^{\dagger}(t)=\begin{pmatrix}\cos\theta(t)e^{i\frac{\alpha(t)}{2}}&-\sin\theta(t)e^{-i\frac{\alpha(t)}{2}}\\ \sin\theta(t)e^{i\frac{\alpha(t)}{2}}&\cos\theta(t)e^{-i\frac{\alpha(t)}{2}}\end{pmatrix}$, (24)
where the parameters $\theta(t)$ and $\alpha(t)$ manipulate the population and the relative phase of both cavity mode $a$ and magnon mode $b$, respectively. From Eqs. (45) and (46), the unitary transformation $\mathcal{V}(t)$ in Eq. (6) which acts as $\mathcal{V}^{\dagger}(t)\mu_{1}(t)\mathcal{V}(t)\rightarrow\mu_{1}(0)$ and $\mathcal{V}^{\dagger}(t)\mu_{2}(t)\mathcal{V}(t)\rightarrow\mu_{2}(0)$, takes the form of
$\mathcal{V}(t)=V_{\alpha}(t)V_{\theta}(t),$ (25)
with
$V_{\alpha}(t)$ $=e^{-i\frac{\alpha(t)-\alpha(0)}{2}(a^{\dagger}a-b^{\dagger}b)},$ (26)
$V_{\theta}(t)$ $=e^{-[\theta(t)-\theta(0)]\left[e^{i\alpha(0)}b^{\dagger}a-e^{-i\alpha(0)}a^{\dagger}b\right]}.$
It is interesting to found that the preceding transformation from the time-dependent ancillary operators to the time-independent ones can also be formally implemented by $V_{\theta}(t)V_{\alpha}(t)]^{\dagger}\mu_{k}(t)V_{\theta}(t)V_{\alpha}(t)\rightarrow\mu_{k}(0)$, $k=1,2$, despite that $V_{\alpha}(t)$ and $V_{\theta}(t)$ are noncommutative, i.e., $V_{\alpha}(t),V_{\theta}(t)]\neq 0$.
Substituting the Hamiltonian (21) and Eqs. (2326) into the upper triangularization condition (10), we obtain the constraints for the coupling strength $J(t)$ and the detuning $\Delta(t)$ as
$J(t)=\Big[\dot{\theta}(t)+\Gamma\cos\alpha(t)\cos 2\theta(t)-\left(\gamma_{a}\cos\varphi_{a}-\gamma_{b}\right)$ (27)
$\times\sin\theta(t)\cos\theta(t)\Big]/\sin[\varphi+\alpha(t)],$
$\Delta(t)=\dot{\alpha}(t)-2\Big[J(t)\cos(\varphi+\alpha(t))\cot 2\theta(t)$
$+\Gamma\frac{\sin\alpha(t)}{\sin 2\theta(t)}+\frac{\gamma_{a}\sin\varphi_{a}}{2}\Big].$
Under Eq. (27), the ancillary operator $\mu_{1}^{\dagger}(t)$ can be activated as a nonadiabatic Heisenberg passage that does not conserve the state-norm. According to Eq. (16), we have
$\mu_{1}^{\dagger}(0)\rightarrow e^{if_{1}(t)}\mu_{1}^{\dagger}(t),$ (28)
where the complex global phase $f_{1}(t)$ can be divided into the real part $f_{r}(t)$ and the imaginary part $f_{i}(t)$ as
$\dot{f}_{1}(t)=\dot{f}_{r}(t)+\dot{f}_{i}(t)$ (29)
with
$\dot{f}_{r}(t)=\frac{1}{2}\Delta(t)\cos 2\theta(t)-J(t)\cos\left[\varphi+\alpha(t)\right]\sin 2\theta(t)$ (30)
$-\frac{\dot{\alpha}(t)}{2}\cos 2\theta(t)+\gamma_{a}\sin\varphi_{a}\cos^{2}\theta(t),$
$\dot{f}_{i}(t)=-i\Big[\gamma_{a}\cos\varphi_{a}\cos^{2}\theta(t)+\gamma_{b}\sin^{2}\theta(t)$
$-\Gamma\cos\alpha\sin 2\theta(t)\Big].$
Equations (23), (24), and (28) indicate that the state evolution of the two-bosonic-mode system is subject to the boundary conditions of $\theta(t)$ and $\alpha(t)$. For example, when $\theta(0)=0$ and $\theta(\tau)=\pi/2$ with $\tau$ the evolution period, $\mu_{1}^{\dagger}(0)=a^{\dagger}\rightarrow\mu_{1}^{\dagger}(\tau)=b^{\dagger}$ and then the initial state of the cavity mode $a$ can be perfectly transferred along the passage $\mu_{1}^{\dagger}(t)$ to the magnon mode $b$ starting from the vacuum state. In addition, the imaginary part of the phase $\dot{f}_{i}(t)$ in Eq. (30) captures the fact that the non-Hermitian component in the Hamiltonian (21) renders the probability nonconservation of the two-mode system during the time evolution. However, in our protocol, the state-norm can be guaranteed to be unit at both beginning and end of the evolution, as long as we have a vanishing integral $\int_{0}^{\tau}\dot{f}_{i}(t)dt=f_{i}(\tau)-f_{i}(0)=0$.
In parallel, for the system dynamics governed by the Hermitian conjugate $H^{\dagger}(t)$ in the dual space, the same conditions in Eq. (27) can activate the ancillary operator $\mu_{2}^{\dagger}(t)$ as the nonadiabatic passage. In particular, the time evolution takes the form of Eq. (19), where the global phase $\dot{f}_{2}(t)=-i[\gamma_{a}\exp(i\varphi_{a})+\gamma_{b}]-\dot{f}_{1}(t)$ with $\dot{f}_{1}(t)$ given by Eqs. (29) and (30). Similar to the passage $\mu_{1}^{\dagger}(t)$, a flexible and perfect state transfer can be completed along the passage $\mu_{2}^{\dagger}(t)$ under appropriate choices of $\theta(t)$ and $\alpha(t)$. The probability conservation can also be ensured in the end of the evolution.
IV State control over cavity and magnon
IV.1 Perfect state transfer
In this section, we use the activated passage $\mu_{1}^{\dagger}(t)$ in Eq. (28) to realize the perfect transfer of arbitrary initial states from the cavity mode to the magnon mode, including the Fock state, the binomial code state (a state of a logical qubit encoding for enhancing noise resilience) Michael et al. (2016), the coherent state, the cat state, and even the thermal state. These transfers are found to be irrelevant to the $\mathcal{PT}$-symmetry of Hamiltonian and the presence or absence of EPs.
We first consider the Hamiltonian in the phase of $\mathcal{PT}$-symmetry. In this case, $\varphi_{a}=\pi$, $\gamma_{a}=\gamma_{b}$, $\Gamma=0$, and $\Delta(t)=0$. Then the constraint on $J(t)$ in Eq. (27) becomes $J(t)=\dot{\theta}(t)/\sin[\varphi+\alpha(t)]$. Suppose the cavity mode and magnon mode initially prepared in the Fock state $|n=5\rangle$ and the vacuum state $|0\rangle$, respectively, i.e., $|\psi(0)\rangle=|5\rangle_{a}|0\rangle_{b}$. Using Eqs. (23), (24), and (28), the system passage can evolve as $\mu_{1}^{\dagger}(0)=a^{\dagger}\rightarrow e^{if_{r}(\tau)}\mu_{1}^{\dagger}(\tau)=e^{if_{r}(\tau)}b^{\dagger}$ under the setting of $\alpha(t)=0$ and
$\theta(t)=\frac{\pi t}{2\tau}.$ (31)
Then the initial state of the two-mode system $|\psi(0)\rangle$
$|\psi(0)\rangle=|5\rangle_{a}|0\rangle_{b}=\frac{(a^{\dagger})^{5}}{\sqrt{5!}}|0\rangle_{a}|0\rangle_{b}=\frac{[\mu_{1}^{\dagger}(0)]^{5}}{\sqrt{5!}}|0\rangle_{a}|0\rangle_{b}$ (32)
is transferred to be
$\frac{[\mu_{1}^{\dagger}(\tau)]^{5}}{\sqrt{5!}}|0\rangle_{a}|0\rangle_{b}=\frac{(b^{\dagger})^{5}}{\sqrt{5!}}|0\rangle_{a}|0\rangle_{b}=|0\rangle_{a}|5\rangle_{b}=|\psi(\tau)\rangle,$ (33)
demonstrating a perfect Fock-state transfer from the cavity mode to the magnon mode.
Under the non-unique condition in Eq. (31), we find that the probability conservation of the system state can be restored at the end of the evolution in the presence or in the absence of exceptional points by appropriately choosing the gain or loss rates and the dissipative coupling strength. Particularly for the Hamiltonian of $\mathcal{PT}$-symmetry, the imaginary part of global phase in Eq. (30) is reduced to be
$\dot{f}_{i}(t)=i\gamma_{a}\cos 2\theta(t).$ (34)
To ensure the vanishing of the time integral over Eq. (34), one can choose the gain or loss rate and dissipative coupling strength as
$\gamma_{a}=\frac{\lambda}{\pi}\dot{\theta}(t),\quad\Gamma=-\frac{\lambda}{2}\dot{\theta}(t),$ (35)
given $\theta(t)$ is set as a linear function of time as in Eq. (31). Here the coefficient $\lambda$ scales the magnitudes of these rates or strength.
The performance of our protocol for state transfer can be evaluated by the fidelity $\mathcal{F}_{\rho}=\langle\psi(t)|\rho|\psi(t)\rangle$, where $|\psi(t)\rangle$ is the pure-state solution of the time-dependent Schrödinger equation (1a) with the Hamiltonian (21). $\rho$ is the density matrix of any interested state, such as initial, intermediate, and target states. For pure states $\rho=|\phi\rangle\langle\phi|$, we have $\mathcal{F}_{\phi}=|\langle\phi|\psi(t)\rangle|^{2}$. When the initial state is a product of Fock states, it is equivalent to show the population dynamics $\mathcal{F}_{n_{1},n_{2}}=|\langle n_{1}|\langle n_{2}|\psi(t)\rangle|^{2}$ during the time evolution.
Figure 2: Fidelity dynamics during the Fock-state transfer $|\psi(0)\rangle=|5\rangle_{a}|0\rangle_{b}\rightarrow|\psi(\tau)\rangle=|0\rangle_{a}|5\rangle_{b}$ under the $\mathcal{PT}$-symmetric Hamiltonian, using the passage $\mu_{1}^{\dagger}(t)$ in the cavity-magnonic system for (a) avoiding EPs and (b) crossing EPs. The associated dynamics of the real and imaginary parts of the energies $E_{\pm}$ in Eq. (22) is plotted in (c) and (d) for avoiding and crossing EPs, respectively. With $\theta(t)$ in Eq. (31), the coherent coupling strength $J(t)$ and the detuning $\Delta(t)$ are constrained by Eq. (27). $\varphi_{a}=\pi$, $\varphi=\pi/2$, $\Gamma=0$, and $\gamma_{a}=\gamma_{b}$, where $\gamma_{a}$ satisfies Eq. (35) with $\lambda=\pi$ in (a) and (c), and $\lambda=4\pi$ in (b) and (d). Then $f_{i}(\tau)-f_{i}(0)=0$ for both avoiding and crossing EPs.
Figure 2 demonstrates perfect nonadiabatic passages $\mu_{1}^{\dagger}(t)$, i.e., $\mathcal{F}_{5,0}(0)=1$ and $\mathcal{F}_{0,5}(\tau)=1$, under a $\mathcal{PT}$-symmetric Hamiltonian. It is found that for both evolutions of (a) avoiding and (b) crossing EPs, the Fock states with conserved excitations other than the initial and target states are scarcely populated. For example, in Fig. 2(a), we have at most $\mathcal{F}_{4,1}(0.19\tau)=0.073$, $\mathcal{F}_{3,2}(0.40\tau)=0.031$, $\mathcal{F}_{2,3}(0.62\tau)=0.032$, and $\mathcal{F}_{1,4}(0.82\tau)=0.074$. The total fidelity $\sum_{n}\mathcal{F}_{n}\equiv\sum_{n_{1},n_{2}}\mathcal{F}_{n_{1},n_{2}}$, being equivalent to the trace of the two-mode system ${\rm Tr}[|\psi(t)\rangle\langle\psi(t)|]$, can represent the probability conservation or non-conservation under a non-Hermitian Hamiltonian. In both Figs. 2(a) and (b), it is found that $\sum_{n}\mathcal{F}_{n}(0<t<\tau)<1$ due to the gain effect of the cavity mode and the loss effect of the magnon mode. Nevertheless, $\sum_{n}\mathcal{F}_{n}(t=\tau)=1$ at the end of the passage. In Figs. 2(c) and (d), we present the dynamics of real and imaginary parts of eigenenergies $E_{\pm}$ due to Eq. (22), relevant to the avoiding and crossing EPs in the parametric setting, respectively. The eigenenergies in Fig. 2(c) verifies that the system dynamics avoids the coalescence of eigenenergies and states. Figure 2(d) indicates that the system crosses EPs when $t=0.18\tau$ and $t=0.82\tau$. It is hardly to find a clear relevance between the perfect state transfer and the existence of the exceptional points of the system spectrum.
Figure 3: Fidelity dynamics during the Fock-state transfer $|\psi(0)\rangle=|5\rangle_{a}|0\rangle_{b}\rightarrow|\psi(\tau)\rangle=|0\rangle_{a}|5\rangle_{b}$ under a $\mathcal{PT}$-symmetric-broken Hamiltonian, using the passage $\mu_{1}^{\dagger}(t)$ in the cavity-magnonic system for (a) avoiding EPs and (b) crossing EPs. The associated dynamics of the real and imaginary parts of the energies $E_{\pm}$ in Eq. (22) is shown in (c) and (d) for avoiding and crossing EPs, respectively. $\varphi_{a}=0$. Both $\Gamma$ and $\gamma_{a}$ are given by Eq. (35) with $\lambda=0.5$ in (a) and (c), and $\lambda=1.2$ in (b) and (d). The other parameters are the same as Fig. 2. And $f_{i}(\tau)-f_{i}(0)=0$ for both avoiding and crossing EPs.
Similar to Figs. 2(a) and (b), Figs. 3(a) and (b) demonstrate the fidelity dynamics $\mathcal{F}_{n_{1},n_{2}}$ of various five-excitation states and their summation $\sum_{n}\mathcal{F}_{n}=\sum_{n_{1},n_{2}}\mathcal{F}_{n_{1},n_{2}}$ within the broken phase of $\mathcal{PT}$-symmetry in the absence and presence of EPs, respectively. Here we set $\varphi_{a}=0$ and $\gamma_{a}=\gamma_{b}$. Then the imaginary part of global phase in Eq. (30) becomes
$\dot{f}_{i}(t)=-i(\gamma_{a}-\Gamma\sin 2\theta).$ (36)
One can check that the gain or loss rate and dissipative coupling strength in Eq. (35) are still applicable to neutralize the imaginary phase under the parametric setting in Eq. (31) for $\theta$.
In both Figs. 3(a) and (b), it is found that the initial population on $|5\rangle_{a}|0\rangle_{0}$ can be completely transferred to $|0\rangle_{a}|5\rangle_{b}$, even when the other five-excitation states could be temporally yet significantly populated during the passage. The Fock states with more number of excitations in the target (magnon) mode are sequently populated until $\mathcal{F}_{0,5}(\tau)=1$ at the end of the time evolution. Note our wavefunction $|\psi(t)\rangle$ is nonunitary and never renormalized. For $0<t<\tau$, the population summation of individual states is not subject to the probability conservation, i.e., $\sum_{n}\mathcal{F}_{n}<1$ or $\sum_{n}\mathcal{F}_{n}>1$, due to the loss effects of both the cavity and magnon modes and the dissipative coupling between them. The peak value of the population summation is found to be $\sum_{n}\mathcal{F}_{n}=1.32$ when $t=0.72\tau$ in Fig. 3(a), around which the dissipative coupling between the cavity and magnon modes dominates the loss effects of both modes. In Fig. 3(a) in the absence of EPs, we have $\mathcal{F}_{4,1}(0.22\tau)=0.35$, $\mathcal{F}_{3,2}(0.34\tau)=0.34$, $\mathcal{F}_{2,3}(0.45\tau)=0.40$, $\mathcal{F}_{1,4}(0.61\tau)=0.55$, and $\mathcal{F}_{0,5}(0.90\tau)=1.07$. And in Fig. 3(b) in the presence of EPs, we have $\mathcal{F}_{4,1}(0.35\tau)=0.24$, $\mathcal{F}_{3,2}(0.50\tau)=0.32$, $\mathcal{F}_{2,3}(0.61\tau)=0.46$, $\mathcal{F}_{1,4}(0.72\tau)=0.74$ and $\mathcal{F}_{0,5}(0.89\tau)=1.40$. Under the conditions in Eq. (35), the state norm becomes unit at the end of the passage. Again the associated evolutions of the real and imaginary parts of the eigenenergies are presented in Figs. 3(c) and (d), respectively. As shown in Fig. 3(c), the eigenenergies do not coalesce during the whole passage, confirming the avoidance of EPs. In Fig. 3(d), an EP occurs at $t=0.27\tau$, also being irrelevant to the system dynamics.
Figure 4: Fidelity dynamics during the perfect transfer of (a) the binomial code state $|\phi_{b},0\rangle\rightarrow|0,\phi_{b}\rangle$ with $|\phi_{b}\rangle=(\sqrt{3}|2\rangle+|6\rangle)/2$ Michael et al. (2016), (b) the coherent state $|\alpha,0\rangle\rightarrow|0,\alpha\rangle$ with $\alpha=5$, (c) the cat state $|\rm cat,0\rangle\rightarrow|0,\rm cat\rangle$, where $|\rm cat\rangle=(|\alpha\rangle+|-\alpha\rangle)/\sqrt{2}$ with $\alpha=5$, and (d) the thermal state $\rho_{\rm th}\otimes|0\rangle\langle 0|\rightarrow|0\rangle\langle 0|\otimes\rho_{\rm th}$, where $\rho_{\rm th}=\sum_{n}p_{n}|n\rangle\langle n|$ with $p_{n}=(\bar{n}^{n})/(1+\bar{n})^{n+1}$ and $\bar{n}=5$. The parameters are the same as Figs. 3(b) and (d).
Far beyond the Fock state, the universal nonadiabatic passage enables perfect transfer of diverse states, including the binomial code state Michael et al. (2016), the coherent state, the cat state, and the thermal state. In Fig. 4(a), the cavity and magnon modes are initially prepared in the binomial code state $|\phi_{b}\rangle=(\sqrt{3}|2\rangle+|6\rangle)/2$ Michael et al. (2016) and the vacuum state $|0\rangle$, respectively. The performance of our protocol can be evaluated by $\mathcal{F}_{\phi_{b},0}=|\langle\phi_{b}|\langle 0|\psi(t)\rangle|^{2}$ and $\mathcal{F}_{0,\phi_{b}}=|\langle 0|\langle\phi_{b}|\psi(t)\rangle|^{2}$. It is found that the binomial code state in the cavity mode can be perfectly transferred to the magnon mode with a unit fidelity. Along the passage, the probability summation during $0<t<\tau$ is not conserved due to the cavity gain and dissipative coupling between the cavity mode and the magnon mode. For example, we have ${\rm Tr}(\rho)=0.631$ when $t=0.234\tau$ and ${\rm Tr}(\rho)=1.63$ when $t=0.782$. However, the probability conservation is restored as ${\rm Tr}(\rho)=1$ when $t=\tau$. Similarly, in Figs. 4(b), (c), and (d) for the cavity mode initially prepared as the coherent state $|\alpha\rangle$ with $\alpha=5$, the cat state $|{\rm cat}\rangle=(|\alpha\rangle+|-\alpha\rangle)/\sqrt{2}$ with $\alpha=5$, and the thermal state $\rho_{\rm th}$ with $\bar{n}=5$, respectively, it is found that all of them can be perfectly transferred to the target mode. Note for the last mixed state, $\mathcal{F}_{{\rm thermal},0}$ and $\mathcal{F}_{0,{\rm thermal}}$ are not complementary to each other in any situation.
In parallel, a quantum state initially prepared in the magnon mode can be perfectly transferred to the cavity mode by activating the passage $\mu_{2}^{\dagger}(t)$ in dual space. It can be simply realized under almost the same conditions as for the passage $\mu_{1}^{\dagger}(t)$, i.e., the parameters employed in Figs. 2, 3, and 4, by replacing $\varphi=\pi/2$ with $\varphi=-\pi/2$.
IV.2 Unidirectional perfect absorber
In this section, we show that the cavity-magnonic system under our UQC theory in Sec. IV.1 can simulate a nonreciprocal behavior under the same control conditions as in Fig. 3 for $\mu_{1}^{\dagger}(t)$ in the ket space. For example, when the Fock state is initially prepared in the magnon mode, i.e., $|\psi(0)\rangle=|0\rangle_{a}|5\rangle_{b}$, the total population of the system is found to be nearly vanishing at the end of time evolution. Note the description about such a phenomenon is meaningful by our theory with no artificial renormalization. Moreover, this unidirectional absorbtion can be interpreted by the coupled-mode theory Xia et al. (2025) about coherent perfect absorber (CPA) Zhang et al. (2017); Chong et al. (2010); Wan et al. (2011); Baranov et al. (2017); Chen et al. (2020); Kim et al. (2025); Xia et al. (2025) to a certain degree.
Under the conditions of Eq. (27) for $\Delta(t)$ with $\varphi_{a}=0$, $\varphi=\pi/2$, $\alpha(t)=0$, and $\gamma_{a}=\gamma_{b}$, the non-Hermitian Hamiltonian in Eq. (21) turns out to be
$H(t)$ $=\begin{pmatrix}a^{\dagger}&b^{\dagger}\end{pmatrix}H^{a}(t)\begin{pmatrix}a\\ b\end{pmatrix}$ (37)
$=\begin{pmatrix}a^{\dagger}&b^{\dagger}\end{pmatrix}\begin{pmatrix}-i\gamma_{a}&i[J(t)+\Gamma]\\ -i[J(t)-\Gamma]&-i\gamma_{a}\end{pmatrix}\begin{pmatrix}a\\ b\end{pmatrix}$.
In the coupled-mode theory Xia et al. (2025), one has to consider the dissipation of the cavity-magnonic system to its external channels Zhang et al. (2017); Chong et al. (2010); Wan et al. (2011); Baranov et al. (2017); Chen et al. (2020); Kim et al. (2025); Xia et al. (2025). This dissipation induces the external loss of the cavity and magnon modes Zhang et al. (2017), by which the total loss rates of the cavity and magnon modes in Eq. (37) can be divided as $\gamma=\gamma_{0}+\gamma_{1}$, where $\gamma_{0}$ and $\gamma_{1}$ represent the intrinsic and external loss rates of the modes, respectively. $\gamma=\gamma_{a}$ or $\gamma_{b}$ in our system.
Following the coupled-mode theory Xia et al. (2025), the system dynamics can be described by the scattering matrix Zhang et al. (2017); Chong et al. (2010); Wan et al. (2011); Baranov et al. (2017); Chen et al. (2020); Kim et al. (2025); Xia et al. (2025). It is defined as
$S(\omega,t)=I-iK^{\dagger}\frac{1}{\omega-H^{a}(t)}K,$ (38)
where $K=\sqrt{2\gamma_{1}}I$ and II is the two-dimensional identity operator. Under the assumption that the incident monochromatic acoustic wave Xia et al. (2025) is resonant with the cavity-magnonic system, i.e., $\omega=0$, the scattering matrix $S$ for $H(t)$ in Eq. (37) can be written as
$S(t)=\begin{pmatrix}S_{11}(t)&S_{12}(t)\\ S_{21}(t)&S_{22}(t)\end{pmatrix}$, (39)
with the reflection coefficients $S_{11}(t)$ and $S_{22}(t)$ and the transmission coefficients $S_{12}(t)$ and $S_{21}(t)$ given by
$S_{11}(t)$ $=S_{22}(t)=1+\frac{2\gamma_{1}\gamma_{a}}{D(t)},$ (40)
$S_{12}(t)$ $=\frac{2\gamma_{1}[J(t)+\Gamma]}{D(t)},\quad S_{21}(t)=-\frac{2\gamma_{1}[J(t)-\Gamma]}{D(t)},$
where $D(t)\equiv-\gamma_{a}^{2}-[J^{2}(t)-\Gamma^{2}]$.
With the same parameters as in Fig. 3, i.e., $J(t)$, $\theta(t)$, and $\gamma_{a}$ and $\Gamma$ are given by Eqs. (27), (31), and (35), respectively, one can find that at the end of running period, $J(\tau)=\dot{\theta}-\Gamma\approx-\Gamma$ when $\lambda\approx 1$. Consequently, the scattering matrix with $t=\tau$ in Eq. (38) becomes
$S(t=\tau)\approx\begin{pmatrix}1-\frac{2\gamma_{1}}{\gamma_{a}}&0\\ S_{21}(\tau)&1-\frac{2\gamma_{1}}{\gamma_{a}}\end{pmatrix}=\begin{pmatrix}0&0\\ S_{21}(\tau)&0\end{pmatrix}$, (41)
where $S_{21}(\tau)\approx 4\gamma_{1}J(\tau)/\gamma_{a}^{2}\neq 0$. Here the second equivalence holds when $\gamma_{1}=\gamma_{a}/2$. Equation (41) indicates that for a nonvacuum initial state in the magnon mode $b$, the total system population can become almost zero at a desired moment, i.e., $a(\tau),b(\tau)]^{T}=S(t=\tau)[0,b(0)]^{T}\approx(0,0)^{T}$. In practice, the non-Hermitian cavity magnonic system now becomes a promising candidate Ramezani et al. (2014); Longhi (2015); Jin et al. (2016); Ramezani et al. (2016); Huang et al. (2017); Jin and Song (2018); Xu and Jin (2024) for the unidirectional perfect absorber under our UQC framework.
Figure 5: Fidelity dynamics about the individual states $|5\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|5\rangle_{b}$ with various coefficients $\lambda$, under a $\mathcal{PT}$-symmetric-broken Hamiltonian. Inset: Numerical results for the logarithm of the nonreciprocity $\log_{10}[S_{21}(\tau)/S_{12}(\tau)]$ as a function of $\lambda$. The other parameters are the same as Fig. 3. The initial state of system is $|\psi(0)\rangle=|0\rangle_{a}|5\rangle_{b}$.
In Fig. 5, we demonstrate the fidelity dynamics of the relevant states $|5\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|5\rangle_{b}$ for various $\lambda$. It is found that while the state $|5\rangle_{a}|0\rangle$ is slightly populated, the population on the state $|0\rangle_{a}|5\rangle_{b}$ decreases monotonically with time, with a decay rate roughly proportional to $\lambda$ for $t\leq 0.3\tau$. In particular, when $t=0.1\tau$, we have $\mathcal{F}_{0,5}=0.713$ for $\lambda=0.5$, $\mathcal{F}_{0,5}=0.601$ for $\lambda=1$ and $\mathcal{F}_{0,5}=0.557$ for $\lambda=1.2$. When $t=0.3\tau$, $\mathcal{F}_{0,5}=0.254$ for $\lambda=0.5$, $\mathcal{F}_{0,5}=0.212$ for $\lambda=1$, and $\mathcal{F}_{0,5}=0.199$ for $\lambda=1.2$. When $t\geq 0.8\tau$, the population of the state $|0\rangle_{a}|5\rangle_{b}$ is almost vanishing for various $\lambda$. This result can be verified by the inset of Fig. 5, where the transition coefficients demonstrate a dramatic non-reciprocal relation $S_{21}(\tau)\gg S_{12}(\tau)$ when $\lambda>0.5$, justifying the unidirectional perfect absorbtion in the cavity-magnonic system. Note for $\lambda=0$, $\log_{10}[S_{21}(\tau)/S_{12}(\tau)]=0$ or $S_{21}(\tau)=S_{12}(\tau)$ implies the disappearance of the non-reciprocal behavior or the bidirectional perfect state transfer in the Hermitian bosonic system Jin and Jing (2025e). It is consistent with the fact that under the condition of Eq. (35) with $\lambda=0$, the non-Hermitian Hamiltonian in Eq. (37) will become a Hermitian one Jin and Jing (2025e), i.e., $H(t)=iJ(t)a^{\dagger}b-iJ(t)b^{\dagger}a$.
V Conclusion
In summary, we propose a versatile method for manipulating the general bosonic system through a non-Hermitian and time-dependent Hamiltonian. In sharp contrast to the conventional treatments of non-Hermitian systems, our theory is built upon the differential manifolds (instantaneous frames) rather than the system’s energy spectra, making the system dynamics independent of $\mathcal{PT}$-symmetry of Hamiltonian and the presence or absence of exceptional points.
In practice, the system dynamics can be described in the ancillary representations. By rotating from the time-dependent to the stationary ancillary representations, a purely geometric gauge potential emerges to shape the time evolution of the system, which is uniquely determined by the formation of ancillary operators. In the stationary representation, the Hamiltonian’s coefficient matrix can be upper triangularized under the triangularization condition with the projection operator, which suffices to activate two Heisenberg passages for flexible and perfect state transfers. Along the transitionless passages, arbitrary initial state can be perfectly transferred between the cavity and magnon modes without artificial normalization, including but not limited to Fock states, coherent states, the superposition of Fock states, the cat states, and even thermal states. The resulting passages do not show a clear relation with both $\mathcal{PT}$-symmetry of Hamiltonian and EPs in the energy spectrum. Moreover, the cavity-magnonic system can be used as a unidirectional perfect absorber under certain conditions. This work essentially extends our universal quantum control to cover the non-Hermitian continuous-variable systems.
Acknowledgments
We acknowledge grant support from the “Pioneer” and “Leading Goose” R&D of Zhejiang Province (Grant No. 2025C01028).
Appendix A A general recipe for ancillary operators
This appendix provides a brief recipe for constructing the ancillary basis operators $\{\mu_{k}(t)\}$ for a general bosonic systems of $N$ modes as well as the $N\times N$ unitary transformation matrix $\mathcal{M}^{\dagger}$ in Eq. (3). Rooted in the geometric structure of the manifold of $\mu_{k}(t)$’s, the adjoint matrix $\mathcal{M}^{\dagger}(t)$ admits the representation Jin and Jing (2025b, e):
$\mathcal{M}^{\dagger}(t)=\left[\vec{M}_{1}(t),\vec{M}_{2}(t),\cdots,\vec{M}_{N}(t)\right]^{T}$ (42)
with the $N$-dimensional row vectors
$\vec{M}_{1}(t)$ $=(\cos\theta_{1}e^{i\frac{\alpha_{1}}{2}},-\sin\theta_{1}e^{-i\frac{\alpha_{1}}{2}},0,\cdots,0),$ (43)
$\vec{M}_{k}(t)$ $=[\cos\theta_{k}e^{i\frac{\alpha_{k}}{2}}\vec{b}_{k-1}(t),-\sin\theta_{k}e^{-i\frac{\alpha_{k}}{2}},0,\cdots,0],$
$\cdots$
$\vec{M}_{N-1}(t)$ $=[\cos\theta_{N-1}e^{i\frac{\alpha_{N-1}}{2}}\vec{b}_{N-2}(t),$
$-\sin\theta_{N-1}e^{-i\frac{\alpha_{N-1}}{2}}],$
$\vec{M}_{N}(t)$ $=\vec{b}_{N-1}(t),$
where $k$ runs from $2$ to N2N-2. Here $\vec{b}_{k}(t)$ is a $k+1$-dimensional bright vector
$\vec{b}_{k}(t)\equiv[\sin\theta_{k}e^{i\frac{\alpha_{k}}{2}}\vec{b}_{k-1}(t),\cos\theta_{k}e^{-i\frac{\alpha_{k}}{2}}],$ (44)
with $1\leq k\leq N-1$ and $\vec{b}_{0}(t)\equiv 1$. In Eqs. (43) and (44), the time-dependence of the parameters $\theta_{k}(t)$ and $\alpha_{k}(t)$ is treated implicitly for clarity. These parameters may be either time-dependent or time-independent.
With the definition in Eqs. (3) and (42), the rotation from the ancillary operators $\{\mu_{k}(t)\}$ to their stationary version $\mathcal{V}^{\dagger}(t)\mu_{k}(t)\mathcal{V}(t)\rightarrow\mu_{k}(0)$ can be performed by Jin and Jing (2025e)
$\mathcal{V}(t)=V_{\alpha_{1}}V_{\theta_{1}}V_{\alpha_{2}}V_{\theta_{2}}\cdots V_{\alpha_{N-1}}V_{\theta_{N-1}}=\prod_{k=1}^{N-1}V_{\alpha_{k}}V_{\theta_{k}},$ (45)
where
$V_{\alpha_{k}}(t)$ $=e^{-i\frac{\delta\alpha_{k}}{2}\left[b_{k-1}^{\dagger}(0)b_{k-1}(0)-a_{k+1}^{\dagger}a_{k+1}\right]},$ (46)
$V_{\theta_{k}}(t)$ $=e^{-\delta\theta_{k}\left[e^{i\alpha_{k}(0)}a_{k+1}^{\dagger}b_{k-1}(0)-e^{-i\alpha_{k}(0)}b_{k-1}^{\dagger}(0)a_{k+1}\right]}$
with $\delta\alpha_{k}=\alpha_{k}(t)-\alpha_{k}(0)$ and $\delta\theta_{k}=\theta_{k}(t)-\theta_{k}(0)$.
Appendix B Derivation of Eq. (21)
This appendix provides a detailed derivation of the non-Hermitian Hamiltonian (21) from the Lindblad master equation. In general, the dynamics of an open cavity magnonic system that is coupled to the traveling waves Metelmann and Clerk (2015); Wang et al. (2019a) can be described by
$\frac{d}{dt}\rho=-i[H_{\rm coh},\rho]+\eta\mathcal{L}[c]\rho+\beta\mathcal{L}[a]\rho+\chi\mathcal{L}[b]\rho,$ (47)
where $H_{\rm coh}=\omega_{a}a^{\dagger}a+\omega_{b}b^{\dagger}b+[J\exp(i\varphi)a^{\dagger}b+{\rm H.c.}]$ represents the eigen-energies of the two modes and the coherent coupling between them. The Lindblad superoperators are defined as $\mathcal{L}[o]\rho=o\rho o^{\dagger}-\{o^{\dagger}o,\rho\}/2$, $o=c,a,b$. The first superoperator is about $c\equiv ua+\exp{(i\Theta)}vb$ with coefficients $u$ and $v$. It implies the cooperative interactions between the two modes and the traveling wave with a damping rate $\eta$. $\Theta=0$ or $\Theta=\pi$ is determined by the propagation direction of the traveling waves. The second and the third superoperators are associated with the individual dissipations of the modes $a$ and $b$, with the damping rates $\beta$ and $\chi$, respectively.
Under Eq. (47), the Schrödinger-picture operator $\mathcal{O}_{S}$ is connected to the Heisenberg-picture operator $\mathcal{O}_{H}(t)$ as
${\rm Tr}\left[\mathcal{O}_{S}\dot{\rho}(t)\right]$ (48)
$={\rm Tr}\left[\mathcal{O}_{S}\left(-i[H_{\rm coh},\rho]+\eta\mathcal{L}[c]\rho+\beta\mathcal{L}[a]\rho+\chi\mathcal{L}[b]\rho\right)\right]$
$={\rm Tr}\left[\left(i[H_{\rm coh},\mathcal{O}_{S}]+\eta\mathcal{L}^{\dagger}[c]\rho+\beta\mathcal{L}^{\dagger}[a]\rho+\chi\mathcal{L}^{\dagger}[b]\right)\rho\right]$
$={\rm Tr}\left[\dot{\mathcal{O}}_{H}(t)\rho(0)\right],$
where the Hermitian conjugate superoperator $\mathcal{L}^{\dagger}[o]\equiv o^{\dagger}\rho o-\{o^{\dagger}o,\rho\}/2$, and the derivation from the second line to the third one has used the cyclic property of the density trace. According to Eq. (48), the time derivative of $\mathcal{O}_{H}(t)$ can be expressed as
$\frac{d}{dt}\mathcal{O}_{H}(t)=i[H_{\rm coh},\mathcal{O}_{S}]+\eta\mathcal{L}^{\dagger}[c]\rho+\beta\mathcal{L}^{\dagger}[a]\rho+\chi\mathcal{L}^{\dagger}[b].$ (49)
Using Eq. (49), the dynamics of the modes $a$ and $b$ can be obtained as
$\frac{d}{dt}a(t)$ $=-i\omega_{a}a-\gamma_{a}a-\left(iJ+e^{i\Theta}\Gamma\right)b,$ (50)
$\frac{d}{dt}b(t)$ $=-i\omega_{b}b-\gamma_{b}b-\left(iJ+e^{i\Theta}\Gamma\right)a,$
under the conditions of $\Gamma=\eta uv$, $\gamma_{a}=\eta v^{2}+\beta$, and $\gamma_{b}=\eta u^{2}+\chi$. The system dynamics described by Eq. (50) is equivalent to that governed by the Hamiltonian (20) with $\varphi_{a}=0$.