### v2

# Universal quantum control over non-Hermitian continuous-variable systems

Abstract
Current studies about the non-Hermitian continuous-variable systems heavily revolved around the singularities in the eigen-spectrum by mimicking their discrete-variable counterparts. The growing discussions over the nonunitary features in the time evolution, however, are limited in scalability and controllability. We here develop a general theory to control an arbitrary number of bosonic modes under the time-dependent non-Hermitian Hamiltonian. Far beyond the subspace of few excitations, our control theory operates in the full Hilbert space within the Heisenberg framework and exploits the gauge potential underlying the instantaneous frames rather than the eigen-spectrum. In particular, the instantaneous frames are defined by time-dependent ancillary operators as linear combinations of the laboratory-frame operators; while the associated gauge potential arises from the unitary transformation connecting the time-dependent and stationary ancillary frames. We find that the upper triangularization condition of the non-Hermitian Hamiltonian’s coefficient matrix in the stationary ancillary frame gives rise to the nonadiabatic passages of two time-dependent ancillary operators, also leading to the exact solutions of the time-dependent Schrödinger equation. At the end of these passages, the probability conservation of the system wavefunction can be automatically restored without artificial normalization. Our theory is exemplified with the perfect and nonreciprocal state transfers in a cavity magnonic system. The former holds for arbitrary initial states and is irrelevant to both parity-time symmetry of the coefficient matrix and exceptional points of the eigen-spectrum; and the latter is consistent with the coherent perfect absorbtion. We essentially constructs the universal quantum control (UQC) theory for the non-Hermitian continuous-variable systems, promising a powerful and reliable approach for their coherent control.
I Introduction
Conventional control protocols for non-Hermitian systems are spectrum-dependent, surrounding such as the broken or unbroken phases of parity-time ($\mathcal{PT}$) symmetry Bender and Boettcher (1998); Ashida et al. (2020); Bender and Hook (2024) of Hamiltonian or scattering matrix Metelmann and Clerk (2015); Ramezani et al. (2014); Longhi (2015); Jin et al. (2016); Ramezani et al. (2016); Huang et al. (2017); Jin and Song (2018); Harder et al. (2018); Grigoryan et al. (2018); Wang et al. (2019a); Xu and Jin (2024) and the presence or absence of the exceptional points (EP) Heiss (2012). $\mathcal{PT}$-symmetry of Hamiltonian provides a relaxed criterion for determining whether the eigen-spectrum of system is real or complex Bender and Boettcher (1998). EPs describe the coalescence of eigenvalues and eigenstates Heiss (2012), at which the biorthogonal condition breaks down Brody (2013). Spectrum-based protocols often impose stringent experimental requirements Ibáñez et al. (2011); Torosov et al. (2013, 2014). More seriously, the extraction of eigen-spectrum becomes increasingly challenging when the quantum system extends from time-independent to time-dependent and from few dimensions to a large scale. The current work aims to establish a universal theory for non-Hermitian continuous-variable systems that does not mimic their discrete-variable counterparts Bender and Boettcher (1998); Ashida et al. (2020); Bender and Hook (2024); Jin and Jing (2025a). It also constitutes a nontrivial extension of the universal quantum control (UQC) Jin and Jing (2025b, c, d, e, f), beyond the closed systems Jin and Jing (2025g) isolated from the external environment.
Non-Hermitian Hamiltonian is the ubiquitous description for continuous-variable systems, such as photonic systems Guo et al. (2009); Regensburger et al. (2012); Ge and Türeci (2013); Peng et al. (2014); Hodaei et al. (2017); Zhang et al. (2020); Bergman et al. (2021), acoustic systems Zhu et al. (2014); Fleury et al. (2015); Christensen et al. (2016); Liu et al. (2018); Wang et al. (2019b); Tang et al. (2020), and cavity magnonic systems Zhang et al. (2017, 2019); Zhao et al. (2020); Wang et al. (2022); Qian et al. (2024); Yang et al. (2024); Zhang et al. (2025). Typically it arises either from the unavoidable energy or material exchange with the environment Metelmann and Clerk (2015); Ramezani et al. (2014); Longhi (2015); Jin et al. (2016); Ramezani et al. (2016); Huang et al. (2017); Jin and Song (2018); Harder et al. (2018); Grigoryan et al. (2018); Wang et al. (2019a); Xu and Jin (2024), or from post-selection over no-quantum-jump trajectories Han et al. (2024), leading to the nonunitary evolution and the violation of probability conservation of the system wavefunction Ashida et al. (2020). Owing to the cooperative coupling to a common reservoir Metelmann and Clerk (2015); Wang et al. (2019a), the non-Hermitian Hamiltonian accounts for the gain or loss effects associated with individual bosonic modes Zhang et al. (2017, 2025) and the dissipative coupling between the system modes Metelmann and Clerk (2015); Harder et al. (2018); Grigoryan et al. (2018); Wang et al. (2019a). In practice, the loss effect on a cavity mode can be effectively converted to the gain effect by applying external microwave fields to its ports Zhang et al. (2017, 2025).
In analog to the two-level systems Ibáñez et al. (2011); Torosov et al. (2013, 2014); Chen et al. (2016); Luan et al. (2022), the shortcut-to-adiabaticity technique was developed to describe the non-Hermitian cavity magnonic systems Zhang et al. (2022) in a highly restrictive subspace. Nevertheless, it is not an approach for the full Hilbert space. Scattering matrix based on the input-output theory Zhang et al. (2017); Ramezani et al. (2014); Longhi (2015); Metelmann and Clerk (2015); Jin et al. (2016); Ramezani et al. (2016); Huang et al. (2017); Jin and Song (2018); Harder et al. (2018); Grigoryan et al. (2018); Wang et al. (2019a); Xu and Jin (2024) formulates a phenomenological and yet rough description for the non-Hermitian continuous-variable systems. Through precise match between coherent and engineered dissipative interaction, this approach can be used to demonstrate the directional quantum amplification Metelmann and Clerk (2015), the unidirectional microwave invisibility Jin and Song (2018); Wang et al. (2019a), and coherent perfect absorption Ramezani et al. (2014); Longhi (2015); Jin et al. (2016); Ramezani et al. (2016); Huang et al. (2017); Jin and Song (2018); Xu and Jin (2024). Most existing works are constrained by the system size and the controllability of target states, excessively depend on the system spectrum, have to repeatedly renormalize the wavefunction during the time evolution, and then cannot be regarded as a systematic and versatile theoretical framework for engineering the non-Hermitian bosonic systems.
In contrast to the eigen-spectrum-based approaches, the universal quantum control framework Jin and Jing (2025b, c, d, e, a, f, g) can provide a broader and more general perspective by exploiting the gauge potential underlying the differential manifold of rotating frames. With the time-dependent Hamiltonian under certain conditions, a UQC theory has been formulated for non-Hermitian discrete-variable systems Jin and Jing (2025a), activating a single nonadiabatic passage in either bra or ket space. In this paper, it is extended to the multiple-mode bosonic systems governed by the non-Hermitian and time-dependent Hamiltonian that is obtained by the Lindblad master equation with the quantum-jump terms retained. Transformation from the time-dependent to the time-independent (stationary) ancillary frames introduces a gauge potential, by which the triangularization condition of the Hamiltonian’s coefficient matrix enables the exact solutions of the Heisenberg equation for two of the ancillary operators. They active two system passages in both bra and ket spaces toward arbitrary target modes. In contrast to the conventional treatments of non-Hermitian systems Daley et al. (2009); Uzdin et al. (2012); Ibáñez and Muga (2014); Ashida and Ueda (2018); Dóra and Moca (2020), our theory is objective, deterministic, and scalable in terms of control target and inherently conserves probability at the end of the passage without artificial normalization. In a cavity magnonic system with gain or loss effects on individual modes and a dissipative coupling between them, it is found that arbitrary state in the cavity mode can be perfectly transferred to the magnon mode and unidirectional perfect absorbtion Ramezani et al. (2014); Longhi (2015); Jin et al. (2016); Ramezani et al. (2016); Huang et al. (2017); Jin and Song (2018); Xu and Jin (2024) can be realized at the operational level if these two modes are exchanged.
The rest of this paper is structured as follows. In Sec. II, we provide a general theory for solving the Heisenberg equation for the bosonic operators driven by a non-Hermitian and time-dependent Hamiltonian. The Hamiltonian’s coefficient matrix can be triangularized in a stationary frame and then yields two ancillary operators to unravel Heisenberg passages. In Sec. III.1, the general theory is applied to the non-Hermitian cavity-magnonic system; and in Sec. III.2, we construct the relevant Heisenberg passages that in general do not conserve the state-norm. Section IV.1 demonstrates that an arbitrary state prepared in the cavity mode can be perfectly transferred to the magnon mode through a designed passage, irrelevant to the parity-time-symmetry and the presence or absence of exceptional points. In addition, Sec. IV.2 shows that under the same setting a unidirectional perfect absorbtion occurs when the two modes are exchanged. The entire work is summarized in Sec. V. Appendix A provides a brief recipe for constructing the ancillary operators as well as the unitary transformation between the time-dependent and stationary ancillary frames. Appendix B details the derivation of the non-Hermitian Hamiltonian based on the Lindblad master equation for the cavity-magnonic system.
II General framework
Consider a general bosonic system consisting of $N$ bosonic modes, associated with the annihilation operators $a_{1}$, $a_{2}$, $\cdots$, $a_{N}$, respectively. The system is controlled by a non-Hermitian Hamiltonian $H(t)$, i.e., $H(t)\neq H^{\dagger}(t)$. Under the assumption of biorthogonal condition Brody (2013), namely that the bra and ket (dual) spaces are equipped with distinct bases, the system dynamics is described by two sets of time-dependent Schrödinger equations as ($\hbar\equiv 1$)
$i\frac{d}{dt}|\psi(t)\rangle$ $=H(t)|\psi(t)\rangle,$ (1a)
$i\frac{d}{dt}|\phi(t)\rangle$ $=H^{\dagger}(t)|\phi(t)\rangle,$ (1b)
where $|\psi(t)\rangle$ and $\langle\phi(t)|$ are the pure-state solutions in the bra and ket spaces, respectively. For the quadratic bosonic systems, the time-dependent Hamiltonian $H(t)$ and its Hermitian-conjugate $H^{\dagger}(t)$ can be expressed as
$H(t)=\vec{a}^{\dagger}H^{a}(t)\vec{a}^{T},\quad H^{\dagger}(t)=\vec{a}^{\dagger}[H^{a}(t)]^{\dagger}\vec{a}^{T},$ (2)
where $\vec{a}\equiv(a_{1},a_{2},\cdots,a_{N})$ and $\vec{a}^{\dagger}=(a_{1}^{\dagger},a_{2}^{\dagger},\cdots,a_{N}^{\dagger})$ are row vectors of operator, and $H^{a}(t)$ is an $N\times N$ time-dependent coefficient matrix with Hermitian-conjugate $H^{a}(t)]^{\dagger}$. $H^{a}(t)\neq[H^{a}(t)]^{\dagger}$. The superscript $T$ denotes the matrix transposition, also transforming a row vector to its corresponding column vector. Due to the noncommunity, non-Hermicity, and the infinite-dimensional Hilbert space of the time-dependent Hamiltonian, solving the Schrödinger equation for non-Hermitian continuous-variable systems is generally challenging. However, our UQC theory Jin and Jing (2025b, c, d, e, a, f, g) can be extended to the non-Hermitian bosonic systems, providing a fundamental framework to partially solve Eqs. (1a) and (1b) in the same time.
Without loss of generality, we start with the system dynamics in the ket space, as given by Eq. (1a) with $H(t)$ in Eq. (2). In the spirit of universal quantum control for Hermitian bosonic systems Jin and Jing (2025g), the system dynamics can be described in the ancillary representation associated with a completed set of time-dependent ancillary operators $\mu_{k}(t)$’s, $1\leq k\leq N$. $\{\mu_{k}(t)\}$ are superposed by the laboratory-frame bosonic operators $\{a_{k}\}$ with an $N\times N$ unitary transformation matrix $\mathcal{M}^{\dagger}(t)$:
$\vec{\mu}_{t}^{T}=\mathcal{M}^{\dagger}(t)\vec{a}^{T},\quad\vec{\mu}_{t}\equiv[\mu_{1}(t),\mu_{2}(t),\cdots,\mu_{N}(t)].$ (3)
A specific unitary transformation matrix $\mathcal{M}^{\dagger}(t)$ and the relevant ancillary operators are constructed in appendix A. Due to Eq. (3) and the unitary feature of $\mathcal{M}^{\dagger}(t)$, one can confirm that the ancillary operators $\mu_{k}(t)$’s satisfy the canonical communication relation, i.e., $\mu_{j}(t),\mu_{k}^{\dagger}(t)]=\delta_{jk}$. Using Eq. (3), the time-dependent Hamiltonian (2) can be formulated as
$H(t)=\vec{\mu}_{t}^{\dagger}H^{\mu}(t)\vec{\mu}_{t}^{T},$ (4)
where
$H^{\mu}(t)=\mathcal{M}^{\dagger}(t)H^{a}(t)\mathcal{M}(t)$ (5)
is the Hamiltonian’s coefficient matrix in the representation of time-dependent ancillary bases $\mu_{k}(t)|{\rm vac}\rangle$ with $|{\rm vac}\rangle$ the vacuum state.
To proceed, we consider the rotation from the time-dependent ancillary operators to the time-independent or stationary ancillary operators, i.e., $\vec{\mu}_{t}\rightarrow\vec{\mu}_{0}$ with $\vec{\mu}_{0}=[\mu_{1}(0),\mu_{2}(0),\cdots,\mu_{N}(0)]$. Such a rotation can be achieved by the unitary transformation $\mathcal{V}(t)$ as $\mathcal{V}^{\dagger}(t)\mu_{k}(t)\mathcal{V}(t)\rightarrow\mu_{k}(0)$, where $\mathcal{V}(t)$ is determined by $\mathcal{M}^{\dagger}(t)$ in Eq. (3) (see appendix A for details). Generally, in the rotating frame with respect to $\mathcal{V}(t)$, we have
$H_{\rm rot}(t)$ $=\mathcal{V}^{\dagger}(t)H(t)\mathcal{V}(t)-i\mathcal{V}^{\dagger}(t)\frac{d\mathcal{V}(t)}{dt}$ (6)
$=\vec{\mu}_{0}^{\dagger}\left[H^{\mu}(t)-\mathcal{A}(t)\right]\vec{\mu}_{0}^{T},$
where the dynamical coefficient matrix $H^{\mu}(t)$ is non-Hermitian, i.e., $H^{\mu}(t)\neq[H^{\mu}(t)]^{\dagger}$. The Hermitian and purely geometric matrix $\mathcal{A}$ represents the gauge potential Kolodrubetz et al. (2017) that is associated with the unitary transformation $\mathcal{V}(t)$. The element of the matrix $\mathcal{A}$ in the $k$th row and $m$th column is defined as $\mathcal{A}_{km}=-i[\mu_{k}^{\dagger}(t),d\mu_{m}(t)/dt]$. In parallel, for the bra-space dynamics in Eq. (1b), the Hamiltonian expressed by the stationary ancillary operators is the Hermitian conjugate of $H_{\rm rot}(t)$, i.e., $H^{\dagger}_{\rm rot}(t)=\vec{\mu}_{0}^{\dagger}[(H^{\mu}(t))^{\dagger}-\mathcal{A}(t)]\vec{\mu}_{0}^{T}$.
Consequently, the time-dependent Schrödinger equations (1) are transformed as
$i\frac{d}{dt}|\psi(t)\rangle_{\rm rot}$ $=H_{\rm rot}(t)|\psi(t)\rangle_{\rm rot},$ (7a)
$i\frac{d}{dt}|\phi(t)\rangle_{\rm rot}$ $=H^{\dagger}_{\rm rot}(t)|\phi(t)\rangle_{\rm rot},$ (7b)
with the rotated pure states
$|\psi(t)\rangle_{\rm rot}=\mathcal{V}^{\dagger}(t)|\psi(t)\rangle,\quad|\phi(t)\rangle_{\rm rot}=\mathcal{V}^{\dagger}(t)|\phi(t)\rangle.$ (8)
The evolution operators for $|\psi(t)\rangle_{\rm rot}$ and $|\phi(t)\rangle_{\rm rot}$ are
$U_{\rm rot}(t)=\hat{T}e^{-i\int_{0}^{t}H_{\rm rot}(s)ds},\quad V_{\rm rot}(t)=\hat{T}e^{-i\int_{0}^{t}H_{\rm rot}^{\dagger}(s)ds},$ (9)
respectively, where $\hat{T}$ is the time-ordering operator.
The preceding rotation to the stationary representation (6) does not directly relieve the difficulty in solving the non-Hermitian Schrödinger equation (7). On one hand, it is not appropriate to apply the commutation condition about the Hamiltonian’s coefficient matrix and the projection operator in the representation of time-independent ancillary modes established for the Hermitian bosonic network Jin and Jing (2025g), since the non-Hermitian coefficient matrix $H^{\mu}(t)-\mathcal{A}(t)$ is generally non-diagonalizable, i.e., $H^{\mu}_{km}(t)\neq H^{\mu}_{mk}(t)$ for $k\neq m$. On the other hand, the triangularization condition about the rotated Hamiltonian for discrete-variable systems Jin and Jing (2025a) also fails here due to the fundamentally distinct statistics of bosonic modes.
Main result.— we prove that the upper triangularization of the coefficient matrix of $H_{\rm rot}(t)$ is a sufficient condition for activating two useful Heisenberg-picture passages, which is only relevant to the differential manifold rather than the spectral characteristics. Specifically, the upper triangularization condition for the coefficient matrix $H^{\mu}(t)-\mathcal{A}(t)$ is given by Jin and Jing (2025a); Ju et al. (2024)
$\left[H^{\mu}(t)-\mathcal{A}(t)\right]\Pi^{k}-\Pi^{k}\left[H^{\mu}(t)-\mathcal{A}(t)\right]^{T}=0,$ (10)
where $\Pi^{k}$, $1\leq k\leq N$, is the projection operator or an $N\times N$ matrix defined by $\Pi^{k}_{jm}=\delta_{jk}\delta_{mk}$.
Proof.— The upper triangularization condition given by Eq. (10) yields the upper triangularized Hamiltonian:
$H_{\rm rot}(t)=\sum_{k=1}^{N}\sum_{m\geq k}^{N}\left[H^{\mu}_{km}(t)-\mathcal{A}_{km}(t)\right]\mu_{k}^{\dagger}(0)\mu_{m}(0).$ (11)
The dynamics of an arbitrary operator $\tilde{\mathcal{O}}_{S}$ can be obtained by $\mathcal{O}_{H}(t)=V_{\rm rot}^{\dagger}(t)\tilde{\mathcal{O}}_{S}U_{\rm rot}(t)$ according to the non-Hermitian Heisenberg equation Miao and Xu (2016) with the evolution operators given by Eq. (9). Under the Hamiltonian (11), the dynamics of the ancillary operators $\mu_{1}^{\dagger}(0)$ and $\mu_{N}(0)$ are found to be decoupled from the others and therefore can be derived as $v_{1}^{\dagger}(t)\equiv V_{\rm rot}^{\dagger}(t)\mu_{1}^{\dagger}(0)U_{\rm rot}(t)$ and $v_{N}(t)\equiv V_{\rm rot}^{\dagger}(t)\mu_{N}(0)U_{\rm rot}(t)$. In particular, we have
$\frac{dv_{1}^{\dagger}(t)}{dt}$ $=iV_{\rm rot}^{\dagger}(t)\left[H_{\rm rot}(t),\mu_{1}^{\dagger}(0)\right]U_{\rm rot}(t)$ (12)
$=i\left[H^{\mu}_{11}(t)-\mathcal{A}_{11}(t)\right]V_{\rm rot}^{\dagger}(t)\mu_{1}^{\dagger}(0)U_{\rm rot}(t)$
$=i\left[H^{\mu}_{11}(t)-\mathcal{A}_{11}(t)\right]v_{1}^{\dagger}(t),$
and
$\frac{dv_{N}(t)}{dt}$ $=iV_{\rm rot}^{\dagger}(t)\left[H_{\rm rot}(t),\mu_{N}(0)\right]U_{\rm rot}(t)$ (13)
$=-i\left[H^{\mu}_{NN}(t)-\mathcal{A}_{NN}(t)\right]V_{\rm rot}^{\dagger}(t)\mu_{N}(0)U_{\rm rot}(t)$
$=-i\left[H^{\mu}_{NN}(t)-\mathcal{A}_{NN}(t)\right]v_{N}(t).$
The solutions to the differential equations (12) and (13) are given by
$v_{1}^{\dagger}(t)=e^{if_{1}(t)}\mu_{1}^{\dagger}(0),\quad v_{N}(t)=e^{-if_{N}(t)}\mu_{N}(0),$ (14)
respectively, where the global phases are defined as
$f_{k}(t)=\int_{0}^{t}ds\left[H^{\mu}_{kk}(s)-\mathcal{A}_{kk}(s)\right],\quad k=1,N.$ (15)
The dynamics of the operators in the picture governed by $H_{\rm rot}(t)$ is related to that in the original picture controlled by $H(t)$ through Eq. (6). According to Eqs. (8) and (14), the dynamics of the ancillary operators $\mu_{1}^{\dagger}(0)$ and $\mu_{N}(0)$ in the original picture can be obtained as
$\mu_{1}^{\dagger}(0)$ $\rightarrow\mathcal{V}(t)v_{1}^{\dagger}(t)\mathcal{V}^{\dagger}(t)$ (16)
$=e^{if_{1}(t)}\mathcal{V}(t)\mu_{1}^{\dagger}(0)\mathcal{V}^{\dagger}(t)=e^{if_{1}(t)}\mu_{1}^{\dagger}(t),$
and
$\mu_{N}(0)$ $\rightarrow\mathcal{V}(t)v_{N}(t)\mathcal{V}^{\dagger}(t)$ (17)
$=e^{-if_{N}(t)}\mathcal{V}(t)\mu_{N}(0)\mathcal{V}^{\dagger}(t)=e^{-if_{N}(t)}\mu_{N}(t).$
Equations (16) and (17) indicate that the ancillary operators $\mu_{1}^{\dagger}(t)$ and $\mu_{N}(t)$ are activated as useful passages toward the desired target mode. The systems initially prepared in the states $F[\mu_{1}^{\dagger}(0)]|{\rm vac}\rangle$ and $F[\mu_{N}(0)]|{\rm vac}\rangle$ will evolve to $F[e^{if_{1}(t)}\mu_{1}^{\dagger}(t)]|{\rm vac}\rangle$ and $F[e^{-if_{N}(t)}\mu_{N}(t)]|{\rm vac}\rangle$ at the time $t$ with the accumulated global phases $f_{1}(t)$ and $f_{N}(t)$, respectively, where $F[\cdot]$ represents an arbitrary function of operators.
In parallel, for the systems governed by the Hermitian-conjugate Hamiltonian $H_{\rm rot}^{\dagger}(t)$ in the dual space, the upper triangularization condition (10) can activate $\mu_{1}(t)$ and $\mu_{N}^{\dagger}(t)$ as nonadiabatic passages in the Heisenberg picture. The Hermitian conjugate of Eq. (11) takes the lower triangularization form of
$H^{\dagger}_{\rm rot}(t)=\sum_{k=1}^{N}\sum_{m\geq k}^{N}\left[(H^{\mu}_{km}(t))^{*}-\mathcal{A}_{km}(t)\right]\mu_{m}^{\dagger}(0)\mu_{k}(0).$ (18)
Similar to the ket-space case, the dynamics of the ancillary operators $\mu_{1}(0)$ and $\mu_{N}^{\dagger}(0)$ can be obtained as
$\mu_{1}(0)\rightarrow e^{-if_{1}^{*}(t)}\mu_{1}(t),\quad\mu_{N}^{\dagger}(0)\rightarrow e^{if_{N}^{*}(t)}\mu_{N}^{\dagger}(t)$ (19)
by using the non-Hermitian Heisenberg equation and the Hamiltonian (18), where the global phase $f_{k}^{*}(t)$, $k=1,N$, is the complex conjugate of $f_{k}(t)$ in Eq. (15).
III Non-Hermitian cavity-magnonic system
III.1 Model and Hamiltonian
Figure 1: Sketch of an open cavity magnonic system under control. More than the coherent exchange coupling with strength $J$ and phase $\varphi$ between the cavity mode $a$ and the magnon mode $b$, the cooperative coupling of the two-mode system to the environment gives rise to the gain or loss rate $\gamma_{a}$ of the cavity mode, the loss rate $\gamma_{b}$ of the magnon mode, and the dissipation coupling $i\Gamma e^{i\Theta}$ between them. $\Theta$ can be 0 or $\pi$ in experiments Wang et al. (2019a).
In this section, we use our UQC theory for the general non-Hermitian bosonic systems in Sec. II to analyze an open two-mode bosonic system that consists of a single-mode cavity and a yttrium iron garnet sphere in the Kittel mode Harder et al. (2018); Grigoryan et al. (2018); Bhoi et al. (2019); Yang et al. (2019); Yao et al. (2019); Yu et al. (2019); Boventer et al. (2020) as shown in Fig. 1. The cavity and magnon modes are represented by the annihilation operators $a$ and $b$, respectively. They are mutually interacted by the time-dependent exchange coupling $J(t)$ with phase $\varphi$. The cooperative coupling between the two-mode system and the environment Metelmann and Clerk (2015); Wang et al. (2019a) leads to the gain or loss rates $\gamma_{a}$ and $\gamma_{b}$ of the cavity mode $a$ and the magnon mode $b$, respectively, as well as a dissipative coupling $i\Gamma e^{i\Theta}$ between them. Then the whole system can be described by a non-Hermitian Hamiltonian Metelmann and Clerk (2015); Wang et al. (2019a):
$H(t)=\left(\omega_{a}-i\gamma_{a}e^{i\varphi_{a}}\right)a^{\dagger}a+\left[\omega_{b}(t)-i\gamma_{b}\right]b^{\dagger}b$ (20)
$+$ $\left[J(t)e^{i\varphi}+i\Gamma e^{i\Theta}\right]a^{\dagger}b+\left[J(t)e^{-i\varphi}+i\Gamma e^{i\Theta}\right]b^{\dagger}a.$
Here $\omega_{a}$ is the free frequency of the cavity mode. $\omega_{b}(t)$ is the time-dependent frequency of the magnon modes, which can be flexibly tuned by the external bias magnetic field $B(t)$ Zhang et al. (2014); Zare Rameshti et al. (2022); Jin and Jing (2024). $\gamma_{b}>0$ is the fixed loss rate of magnon mode. $\gamma_{a}>0$ represents the gain or loss effect on the cavity mode when $\varphi_{a}=\pi$ or $\varphi_{a}=0$, respectively Zhang et al. (2017, 2025). The phase $\Theta=0$ or $\Theta=\pi$ of the dissipative coupling is determined by the propagation direction of the traveling waves Wang et al. (2019a). For simplicity, we set $\Theta=0$ in the following. The details about the derivation of the non-Hermitian Hamiltonian (20) through the Lindblad master equation can be found in appendix B with $\varphi_{a}=0$, despite both $\varphi_{a}$ and $\Gamma$ can have many choices Wang et al. (2019a); Zhang et al. (2017, 2025).
In the rotating frame with respect to $H_{0}(t)=\omega_{0}(t)(a^{\dagger}a+b^{\dagger}b)$, the full Hamiltonian (20) is transformed as
$H(t)$ $=\left[\frac{\Delta(t)}{2}-i\gamma_{a}e^{i\varphi_{a}}\right]a^{\dagger}a-\left[\frac{\Delta(t)}{2}+i\gamma_{b}\right]b^{\dagger}b$ (21)
$+\left[J(t)e^{i\varphi}+i\Gamma\right]a^{\dagger}b+\left[J(t)e^{-i\varphi}+i\Gamma\right]b^{\dagger}a,$
where the detuning $\Delta(t)$ follows $\Delta(t)/2=\omega_{a}-\omega_{0}(t)=-\omega_{b}(t)+\omega_{0}(t)$. One can find that the system Hamiltonian (21) satisfies $\mathcal{PT}$-symmetry Ashida et al. (2020); Bender and Hook (2024) under the setting of $\varphi_{a}=\pi$, $\gamma_{a}=\gamma_{b}$, $\Delta(t)=0$, and the vanishing dissipative coupling $\Gamma=0$. In fact, the Hamiltonian eigenenergies
$E_{\pm}=-i\frac{\gamma_{a}e^{i\varphi_{a}}+\gamma_{b}}{2}\pm\frac{1}{2}\Big\{\Delta^{2}-\left(\gamma_{a}e^{i\varphi_{a}}-\gamma_{b}\right)^{2}$ (22)
$+4(J^{2}-\Gamma^{2})+i\left[8J\Gamma\cos\varphi-2\Delta\left(\gamma_{a}e^{i\varphi_{a}}-\gamma_{b}\right)\right]\Big\}^{1/2}.$
become real-valued, i.e., $E_{\pm}=\pm\sqrt{J^{2}-\gamma_{a}^{2}}$, when $H(t),\mathcal{PT}]=0$ and $J>\gamma_{a}$. Otherwise, they are complex and the Hamiltonian is in the $\mathcal{PT}$-symmetry-broken phase. Whether holding $\mathcal{PT}$-symmetry or not, the system can approach EP under the conditions of $\Delta(t)=0$, $\varphi=\pi/2$, $\gamma_{a}=\gamma_{b}$, and $J=\pm\Gamma$ (when $\varphi_{a}=0$) or $J=\pm\sqrt{\Gamma^{2}+\gamma_{a}^{2}}$ (when $\varphi_{a}=\pi$).
Conventionally, the control protocols proposed for non-Hermitian systems are based on their spectrum characteristics Ashida et al. (2020); Bender and Hook (2024), focusing on $\mathcal{PT}$-symmetry and EPs. In contrast, our UQC theory in Sec. II is essentially proposed in the instantaneous frames that are irrespective to the spectral properties. In the following section, we will demonstrate that an arbitrary state can be perfectly transferred from the cavity mode to magnon mode, regardless of whether the system in the unbroken or broken phases of $\mathcal{PT}$-symmetry and whether EPs are present or not.
III.2 Universal Heisenberg passages of two operators
With Eq. (21), we construct the universal passages for the non-Hermitian cavity-magnonic system, which give rise to nonunitary evolution of the system state and the final state can be automatically normalized. According to Eqs. (42) and (43), the time-dependent ancillary operators of an arbitrary two-mode system can be written as
$\mu_{1}(t),\mu_{2}(t)]^{T}=\mathcal{M}^{\dagger}(t)(a,b)^{T}$ (23)
with a $2\times 2$ unitary transformation matrix
$\mathcal{M}^{\dagger}(t)=\begin{pmatrix}\cos\theta(t)e^{i\frac{\alpha(t)}{2}}&-\sin\theta(t)e^{-i\frac{\alpha(t)}{2}}\\ \sin\theta(t)e^{i\frac{\alpha(t)}{2}}&\cos\theta(t)e^{-i\frac{\alpha(t)}{2}}\end{pmatrix}$, (24)
where the parameters $\theta(t)$ and $\alpha(t)$ manipulate the population and the relative phase of both cavity mode $a$ and magnon mode $b$, respectively. From Eqs. (45) and (46), the unitary transformation $\mathcal{V}(t)$ in Eq. (6) which acts as $\mathcal{V}^{\dagger}(t)\mu_{1}(t)\mathcal{V}(t)\rightarrow\mu_{1}(0)$ and $\mathcal{V}^{\dagger}(t)\mu_{2}(t)\mathcal{V}(t)\rightarrow\mu_{2}(0)$, can take the form of
$\mathcal{V}(t)=V_{\alpha}(t)V_{\theta}(t),$ (25)
with
$V_{\alpha}(t)$ $=e^{-i\frac{\alpha(t)-\alpha(0)}{2}(a^{\dagger}a-b^{\dagger}b)},$ (26)
$V_{\theta}(t)$ $=e^{-[\theta(t)-\theta(0)]\left[e^{i\alpha(0)}b^{\dagger}a-e^{-i\alpha(0)}a^{\dagger}b\right]}.$
It is interesting to find that the preceding transformation from the time-dependent ancillary operators to the time-independent ones can also be formally implemented by $V_{\theta}(t)V_{\alpha}(t)]^{\dagger}\mu_{k}(t)V_{\theta}(t)V_{\alpha}(t)\rightarrow\mu_{k}(0)$, $k=1,2$, despite that $V_{\alpha}(t),V_{\theta}(t)]\neq 0$.
Substituting the Hamiltonian (21) and Eqs. (2326) into the upper triangularization condition (10), we obtain the constraints for the coupling strength $J(t)$ and the detuning $\Delta(t)$ as
$J(t)=\Big[\dot{\theta}(t)+\Gamma\cos\alpha(t)\cos 2\theta(t)-\left(\gamma_{a}\cos\varphi_{a}-\gamma_{b}\right)$ (27)
$\times\sin\theta(t)\cos\theta(t)\Big]/\sin[\varphi+\alpha(t)],$
$\Delta(t)=\dot{\alpha}(t)-2\Big[J(t)\cos(\varphi+\alpha(t))\cot 2\theta(t)$
$+\Gamma\frac{\sin\alpha(t)}{\sin 2\theta(t)}+\frac{\gamma_{a}\sin\varphi_{a}}{2}\Big].$
Under Eq. (27), the ancillary operator $\mu_{1}^{\dagger}(t)$ can be activated as a nonadiabatic Heisenberg passage that does not conserve the state-norm. According to Eq. (16), we have
$\mu_{1}^{\dagger}(0)\rightarrow e^{if_{1}(t)}\mu_{1}^{\dagger}(t),$ (28)
where the complex global phase $f_{1}(t)$ can be divided into the real part $f_{r}(t)$ and the imaginary part $f_{i}(t)$ as
$\dot{f}_{1}(t)=\dot{f}_{r}(t)+\dot{f}_{i}(t)$ (29)
with
$\dot{f}_{r}(t)=\frac{1}{2}\Delta(t)\cos 2\theta(t)-J(t)\cos\left[\varphi+\alpha(t)\right]\sin 2\theta(t)$ (30)
$-\frac{\dot{\alpha}(t)}{2}\cos 2\theta(t)+\gamma_{a}\sin\varphi_{a}\cos^{2}\theta(t),$
$\dot{f}_{i}(t)=-i\Big[\gamma_{a}\cos\varphi_{a}\cos^{2}\theta(t)+\gamma_{b}\sin^{2}\theta(t)$
$-\Gamma\cos\alpha\sin 2\theta(t)\Big].$
Equations (23), (24), and (28) indicate that the state evolution of the two-bosonic-mode system is subject to the boundary conditions of $\theta(t)$ and $\alpha(t)$. For example, when $\theta(0)=0$ and $\theta(\tau)=\pi/2$ with $\tau$ the evolution period, $\mu_{1}^{\dagger}(0)=a^{\dagger}\rightarrow\mu_{1}^{\dagger}(\tau)=b^{\dagger}$ and then the initial state of the cavity mode $a$ can be perfectly transferred along the passage $\mu_{1}^{\dagger}(t)$ to the magnon mode $b$ starting from the vacuum state. In addition, the imaginary part of the phase $\dot{f}_{i}(t)$ in Eq. (30) captures the fact that the non-Hermitian component in the Hamiltonian (21) renders the probability nonconservation of the two-mode system during the time evolution. However, in our protocol, the state-norm can be guaranteed to be unit at both beginning and end of the evolution, as long as we have a vanishing integral $\int_{0}^{\tau}\dot{f}_{i}(t)dt=f_{i}(\tau)-f_{i}(0)=0$.
In parallel, for the system dynamics governed by the Hermitian conjugate $H^{\dagger}(t)$ in the dual space, the same conditions in Eq. (27) can activate the ancillary operator $\mu_{2}^{\dagger}(t)$ as the nonadiabatic passage. In particular, the time evolution takes the form of Eq. (19), where the global phase $\dot{f}_{2}^{*}(t)=i[\gamma_{a}\exp(-i\varphi_{a})+\gamma_{b}]-\dot{f}_{1}^{*}(t)$ with $\dot{f}_{1}^{*}(t)$ the complex conjugate of $\dot{f}_{1}(t)$ given by Eqs. (29) and (30). Similar to $\mu_{1}^{\dagger}(t)$, a flexible and perfect state transfer can be implemented along the passage $\mu_{2}^{\dagger}(t)$ under appropriate choices of $\theta(t)$ and $\alpha(t)$. The probability conservation can also be ensured in the end of the evolution.
IV State control over cavity and magnon
IV.1 Perfect state transfer
In this section, we use the activated passage $\mu_{1}^{\dagger}(t)$ in Eq. (28) to realize the perfect transfer of arbitrary initial states from the cavity mode to the magnon mode, including the Fock state, the binomial code state (a state of a logical qubit encoding for enhancing noise resilience) Michael et al. (2016), the coherent state, the cat state, and even the thermal state. These transfers are found to be irrelevant to the $\mathcal{PT}$-symmetry of Hamiltonian and the presence or absence of EPs.
We first consider the Hamiltonian in the unbroken phase of $\mathcal{PT}$-symmetry. In this case, $\varphi_{a}=\pi$, $\gamma_{a}=\gamma_{b}$, $\Gamma=0$, and $\Delta(t)=0$. Consequently, the constraint on $J(t)$ in Eq. (27) becomes $J(t)=\dot{\theta}(t)/\sin[\varphi+\alpha(t)]$. Using Eqs. (23), (24), and (28), the system passage can evolve as $\mu_{1}^{\dagger}(0)=a^{\dagger}\rightarrow e^{if_{r}(\tau)}\mu_{1}^{\dagger}(\tau)=e^{if_{r}(\tau)}b^{\dagger}$ under the setting of $\alpha(t)=0$ and
$\theta(t)=\frac{\pi t}{2\tau}.$ (31)
Assume that the cavity mode and magnon mode are initially prepared in the Fock state $|n=5\rangle$ and the vacuum state $|0\rangle$, respectively, i.e.,
$|\psi(0)\rangle=|5\rangle_{a}|0\rangle_{b}=\frac{(a^{\dagger})^{5}}{\sqrt{5!}}|0\rangle_{a}|0\rangle_{b}=\frac{[\mu_{1}^{\dagger}(0)]^{5}}{\sqrt{5!}}|0\rangle_{a}|0\rangle_{b}.$ (32)
It will be finally transformed to be
$\frac{[\mu_{1}^{\dagger}(\tau)]^{5}}{\sqrt{5!}}|0\rangle_{a}|0\rangle_{b}=\frac{(b^{\dagger})^{5}}{\sqrt{5!}}|0\rangle_{a}|0\rangle_{b}=|0\rangle_{a}|5\rangle_{b}=|\psi(\tau)\rangle$ (33)
up to an irrelevant phase, demonstrating a perfect Fock-state transfer from the cavity mode to the magnon mode.
Under the non-unique condition in Eq. (31), we find that the probability conservation of the system state can be restored at the end of the evolution in the presence or in the absence of exceptional points by appropriately choosing the gain or loss rates and the dissipative coupling strength. For the Hamiltonian of $\mathcal{PT}$-symmetry, the imaginary part of the global phase in Eq. (30) is reduced to be
$\dot{f}_{i}(t)=i\gamma_{a}\cos 2\theta(t).$ (34)
To ensure the vanishing of the time integral over $\dot{f}_{i}(t)$ in Eq. (34), one can choose the gain or loss rate and dissipative coupling strength as
$\gamma_{a}=\frac{\lambda}{\pi}\dot{\theta}(t),\quad\Gamma=-\frac{\lambda}{2}\dot{\theta}(t),$ (35)
given $\theta(t)$ a linear function of time in Eq. (31). Here the positive factor $\lambda$ scales the magnitudes of these rates or strength.
The performance of our protocol for state transfer can be evaluated by the fidelity $\mathcal{F}_{\rho}=\langle\psi(t)|\rho|\psi(t)\rangle$, where $|\psi(t)\rangle$ is the pure-state solution of the time-dependent Schrödinger equation (1a) with the Hamiltonian (21). $\rho$ is the density matrix of any interested state, such as initial, intermediate, and target states. For pure states $\rho=|\phi\rangle\langle\phi|$, we have $\mathcal{F}_{\phi}=|\langle\phi|\psi(t)\rangle|^{2}$. When the initial state is a product of Fock states, it is equivalent to show the population dynamics $\mathcal{F}_{n_{1},n_{2}}=|\langle n_{1}|\langle n_{2}|\psi(t)\rangle|^{2}$ during the time evolution.
Figure 2: Fidelity dynamics during the Fock-state transfer $|\psi(0)\rangle=|5\rangle_{a}|0\rangle_{b}\rightarrow|\psi(\tau)\rangle=|0\rangle_{a}|5\rangle_{b}$ under the $\mathcal{PT}$-symmetric Hamiltonian, using the passage $\mu_{1}^{\dagger}(t)$ in the cavity-magnonic system for (a) avoiding EPs and (b) crossing EPs. The associated dynamics of the real and imaginary parts of the energies $E_{\pm}$ in Eq. (22) is plotted in (c) and (d) for avoiding and crossing EPs, respectively. With $\theta(t)$ in Eq. (31), the coherent coupling strength $J(t)$ and the detuning $\Delta(t)$ are constrained by Eq. (27). $\varphi_{a}=\pi$, $\varphi=\pi/2$, $\Gamma=0$, and $\gamma_{a}=\gamma_{b}$, where $\gamma_{a}$ satisfies Eq. (35) with $\lambda=\pi$ in (a) and (c), and $\lambda=4\pi$ in (b) and (d). Then $f_{i}(\tau)-f_{i}(0)=0$ for both avoiding and crossing EPs.
Figure 2 demonstrates perfect nonadiabatic passage $\mu_{1}^{\dagger}(t)$, i.e., $\mathcal{F}_{5,0}(0)=1$ and $\mathcal{F}_{0,5}(\tau)=1$, under a $\mathcal{PT}$-symmetric Hamiltonian. It is found that for both evolutions of (a) avoiding and (b) crossing EPs, the Fock states with conserved excitations other than the initial and target states are scarcely populated. For example, in Fig. 2(a), we have at most $\mathcal{F}_{4,1}(0.19\tau)=0.073$, $\mathcal{F}_{3,2}(0.40\tau)=0.031$, $\mathcal{F}_{2,3}(0.62\tau)=0.032$, and $\mathcal{F}_{1,4}(0.82\tau)=0.074$. The total fidelity $\sum_{n}\mathcal{F}_{n}\equiv\sum_{n_{1},n_{2}}\mathcal{F}_{n_{1},n_{2}}$, being equivalent to the trace of the two-mode system ${\rm Tr}[|\psi(t)\rangle\langle\psi(t)|]$, can represent the probability conservation or non-conservation under a non-Hermitian Hamiltonian. In both Figs. 2(a) and (b), it is found that $\sum_{n}\mathcal{F}_{n}(0<t<\tau)<1$ due to the gain effect of the cavity mode and the loss effect of the magnon mode. Nevertheless, $\sum_{n}\mathcal{F}_{n}(t=\tau)$ always converges to unit at the end of the passage. In Figs. 2(c) and (d), we present the dynamics of real and imaginary parts of eigenenergies $E_{\pm}$ due to Eq. (22), relevant to the avoiding and crossing EPs in the parametric setting, respectively. The curves in Fig. 2(c) verify that the system dynamics avoids the coalescence of eigenenergies and eigenstates. Figure 2(d) indicates that the system crosses EPs when $t=0.18\tau$ and $t=0.82\tau$. It is hardly to find a clear connection between the perfect state transfer and the existence of the exceptional points of the system spectrum.
Figure 3: Fidelity dynamics during the Fock-state transfer $|\psi(0)\rangle=|5\rangle_{a}|0\rangle_{b}\rightarrow|\psi(\tau)\rangle=|0\rangle_{a}|5\rangle_{b}$ under a $\mathcal{PT}$-symmetric-broken Hamiltonian, using the passage $\mu_{1}^{\dagger}(t)$ in the cavity-magnonic system for (a) avoiding EPs and (b) crossing EPs. The associated dynamics of the real and imaginary parts of the energies $E_{\pm}$ in Eq. (22) is shown in (c) and (d) for avoiding and crossing EPs, respectively. $\varphi_{a}=0$. Both $\Gamma$ and $\gamma_{a}$ are given by Eq. (35) with $\lambda=0.5$ in (a) and (c), and $\lambda=1.2$ in (b) and (d). The other parameters are the same as Fig. 2. And $f_{i}(\tau)-f_{i}(0)=0$ for both avoiding and crossing EPs.
Similar to Figs. 2(a) and (b), Figs. 3(a) and (b) demonstrate the fidelity dynamics $\mathcal{F}_{n_{1},n_{2}}$ of various five-excitation product-states and their summation $\sum_{n}\mathcal{F}_{n}=\sum_{n_{1},n_{2}}\mathcal{F}_{n_{1},n_{2}}$ within the broken phase of $\mathcal{PT}$-symmetry in the absence and presence of EPs, respectively. Here we set $\varphi_{a}=0$ and $\gamma_{a}=\gamma_{b}$. Then the imaginary part of the global phase in Eq. (30) becomes
$\dot{f}_{i}(t)=-i(\gamma_{a}-\Gamma\sin 2\theta).$ (36)
One can check that the gain or loss rate and dissipative coupling strength in Eq. (35) are still applicable to neutralize the imaginary phase under the parametric setting in Eq. (31) for $\theta(t)$.
In both Figs. 3(a) and (b), it is found that the initial population on $|5\rangle_{a}|0\rangle_{0}$ can be completely transferred to $|0\rangle_{a}|5\rangle_{b}$, even when the other five-excitation states could be temporally yet significantly populated during the passage. The Fock states with more excitations of the target (magnon) mode are sequently populated until $\mathcal{F}_{0,5}(\tau)=1$ at the end of the time evolution. For $0<t<\tau$, the summation of individual populations is not conserved, i.e., $\sum_{n}\mathcal{F}_{n}<1$ or $\sum_{n}\mathcal{F}_{n}>1$, due to the loss effects of both the cavity and magnon modes and the dissipative coupling between them. The peak value of the population summation is found to be $\sum_{n}\mathcal{F}_{n}=1.32$ when $t=0.72\tau$ in Fig. 3(a), around which the dissipative coupling between the cavity and magnon modes dominates the loss of both modes. In Fig. 3(a) with no EPs, we have $\mathcal{F}_{4,1}(0.22\tau)=0.35$, $\mathcal{F}_{3,2}(0.34\tau)=0.34$, $\mathcal{F}_{2,3}(0.45\tau)=0.40$, $\mathcal{F}_{1,4}(0.61\tau)=0.55$, and $\mathcal{F}_{0,5}(0.90\tau)=1.07$. And in Fig. 3(b) with EPs, we have $\mathcal{F}_{4,1}(0.35\tau)=0.24$, $\mathcal{F}_{3,2}(0.50\tau)=0.32$, $\mathcal{F}_{2,3}(0.61\tau)=0.46$, $\mathcal{F}_{1,4}(0.72\tau)=0.74$, and $\mathcal{F}_{0,5}(0.89\tau)=1.40$. Under the conditions in Eq. (35), the state norm becomes unit at the end of the passage. Again the associated evolutions of the real and imaginary parts of the eigenenergies are presented in Figs. 3(c) and (d), respectively. As shown in Fig. 3(c), the eigenenergies do not coalesce during the whole passage, confirming the avoidance of EPs. In Fig. 3(d), an EP occurs at $t=0.27\tau$, also being irrelevant to the system dynamics.
Figure 4: Fidelity dynamics during the perfect transfer of (a) the binomial code state $|\phi_{b},0\rangle\rightarrow|0,\phi_{b}\rangle$ with $|\phi_{b}\rangle=(\sqrt{3}|2\rangle+|6\rangle)/2$ Michael et al. (2016), (b) the coherent state $|\alpha,0\rangle\rightarrow|0,\alpha\rangle$ with $\alpha=5$, (c) the cat state $|\rm cat,0\rangle\rightarrow|0,\rm cat\rangle$, where $|\rm cat\rangle=(|\alpha\rangle+|-\alpha\rangle)/\sqrt{2}$ with $\alpha=5$, and (d) the thermal state $\rho_{\rm th}\otimes|0\rangle\langle 0|\rightarrow|0\rangle\langle 0|\otimes\rho_{\rm th}$, where $\rho_{\rm th}=\sum_{n}p_{n}|n\rangle\langle n|$ with $p_{n}=(\bar{n}^{n})/(1+\bar{n})^{n+1}$ and $\bar{n}=5$. The parameters are the same as Figs. 3(b) and (d).
Far beyond the Fock state, the universal Heisenberg passage can unravel perfect transfer of diverse states. In Fig. 4(a), the cavity and magnon modes are initially prepared in the binomial code state $|\phi_{b}\rangle=(\sqrt{3}|2\rangle+|6\rangle)/2$ Michael et al. (2016) and the vacuum state $|0\rangle$, respectively. The performance of our protocol can be evaluated by $\mathcal{F}_{\phi_{b},0}=|\langle\phi_{b}|\langle 0|\psi(t)\rangle|^{2}$ and $\mathcal{F}_{0,\phi_{b}}=|\langle 0|\langle\phi_{b}|\psi(t)\rangle|^{2}$. It is found that the binomial code state in the cavity mode can be perfectly transferred to the magnon mode with a unit fidelity. Along the passage, the non-Hermitian system during $0<t<\tau$ is not subject to the probability conservation. For example, we have ${\rm Tr}(\rho)=0.631$ when $t=0.234\tau$ and ${\rm Tr}(\rho)=1.63$ when $t=0.782$. However, the probability conservation is restored as ${\rm Tr}(\rho)=1$ when $t=\tau$. Similarly, in Figs. 4(b), (c), and (d) for the cavity mode initially prepared as the coherent state $|\alpha\rangle$ with $\alpha=5$, the cat state $|{\rm cat}\rangle=(|\alpha\rangle+|-\alpha\rangle)/\sqrt{2}$ with $\alpha=5$, and the thermal state $\rho_{\rm th}$ with $\bar{n}=5$, respectively, it is found that all of them can be perfectly transferred to the target mode. Note for the thermal state, $\mathcal{F}_{{\rm thermal},0}$ and $\mathcal{F}_{0,{\rm thermal}}$ are not complementary to each other in any situation.
In parallel, a quantum state initially prepared in the magnon mode can be perfectly transferred to the cavity mode by activating the passage $\mu_{2}^{\dagger}(t)$ in the dual space. It can be realized under almost the same conditions as for the passage $\mu_{1}^{\dagger}(t)$, except that $\varphi=\pi/2$ is replaced with $\varphi=-\pi/2$.
IV.2 Unidirectional perfect absorber
In this section, we show that the cavity-magnonic system under our UQC theory in Sec. IV.1 can simulate a nonreciprocal behavior under the same control conditions as in Fig. 3 for $\mu_{1}^{\dagger}(t)$ in the ket space. For example, when the Fock state is initially prepared in the magnon mode, i.e., $|\psi(0)\rangle=|0\rangle_{a}|5\rangle_{b}$, the total population of the system is found to be nearly vanishing at the end of time evolution. Note the description about such a phenomenon is meaningful by our theory with no artificial renormalization. To a certain degree, this unidirectional absorbtion can be interpreted by the coupled-mode theory Xia et al. (2025) about coherent perfect absorber (CPA) Zhang et al. (2017); Chong et al. (2010); Wan et al. (2011); Baranov et al. (2017); Chen et al. (2020); Kim et al. (2025); Xia et al. (2025).
Under the conditions of Eq. (27) for $\Delta(t)$ with $\varphi_{a}=0$, $\varphi=\pi/2$, $\alpha(t)=0$, and $\gamma_{a}=\gamma_{b}$, the non-Hermitian Hamiltonian in Eq. (21) turns out to be
$H(t)$ $=\begin{pmatrix}a^{\dagger}&b^{\dagger}\end{pmatrix}H^{a}(t)\begin{pmatrix}a\\ b\end{pmatrix}$ (37)
$=\begin{pmatrix}a^{\dagger}&b^{\dagger}\end{pmatrix}\begin{pmatrix}-i\gamma_{a}&iJ(t)+i\Gamma\\ -iJ(t)+i\Gamma&-i\gamma_{a}\end{pmatrix}\begin{pmatrix}a\\ b\end{pmatrix}$.
In the coupled-mode theory Xia et al. (2025), one has to consider the dissipation of the cavity-magnonic system to its external channels Zhang et al. (2017); Chong et al. (2010); Wan et al. (2011); Baranov et al. (2017); Chen et al. (2020); Kim et al. (2025); Xia et al. (2025). This dissipation induces the external loss of the cavity and magnon modes Zhang et al. (2017), by which the total loss rates of the cavity and magnon modes in Eq. (37) can be divided as $\gamma=\gamma_{0}+\gamma_{1}$, where $\gamma_{0}$ and $\gamma_{1}$ represent the intrinsic and external loss rates of the modes, respectively. $\gamma=\gamma_{a}$ or $\gamma_{b}$ in our system.
Following the coupled-mode theory Xia et al. (2025), the system dynamics can be described by the scattering matrix Zhang et al. (2017); Chong et al. (2010); Wan et al. (2011); Baranov et al. (2017); Chen et al. (2020); Kim et al. (2025); Xia et al. (2025). It is defined as
$S(\omega,t)=I-iK^{\dagger}\frac{1}{\omega-H^{a}(t)}K,$ (38)
where $K=\sqrt{2\gamma_{1}}I$ and II is the two-dimensional identity operator. Under the assumption that the incident monochromatic acoustic wave Xia et al. (2025) is resonant with the cavity-magnonic system, i.e., $\omega=0$, the scattering matrix $S$ for $H(t)$ in Eq. (37) can be written as
$S(t)=\begin{pmatrix}S_{11}(t)&S_{12}(t)\\ S_{21}(t)&S_{22}(t)\end{pmatrix}$, (39)
with the reflection coefficients $S_{11}(t)$ and $S_{22}(t)$ and the transmission coefficients $S_{12}(t)$ and $S_{21}(t)$ given by
$S_{11}(t)$ $=S_{22}(t)=1+\frac{2\gamma_{1}\gamma_{a}}{D(t)},$ (40)
$S_{12}(t)$ $=\frac{2\gamma_{1}[J(t)+\Gamma]}{D(t)},\quad S_{21}(t)=-\frac{2\gamma_{1}[J(t)-\Gamma]}{D(t)},$
where $D(t)\equiv-\gamma_{a}^{2}-[J^{2}(t)-\Gamma^{2}]$.
With the same parameters as in Fig. 3, i.e., $J(t)$, $\theta(t)$, and $\gamma_{a}$ and $\Gamma$ are given by Eqs. (27), (31), and (35), respectively, one can find that at the end of running period, $J(\tau)=\dot{\theta}-\Gamma\approx-\Gamma$ when $\lambda\approx 1$. Consequently, the scattering matrix with $t=\tau$ in Eq. (38) becomes
$S(t=\tau)\approx\begin{pmatrix}1-\frac{2\gamma_{1}}{\gamma_{a}}&0\\ S_{21}(\tau)&1-\frac{2\gamma_{1}}{\gamma_{a}}\end{pmatrix}=\begin{pmatrix}0&0\\ S_{21}(\tau)&0\end{pmatrix}$, (41)
where $S_{21}(\tau)\approx 4\gamma_{1}J(\tau)/\gamma_{a}^{2}\neq 0$. Here the second equivalence holds when $\gamma_{1}=\gamma_{a}/2$. Equation (41) indicates that for a nonvacuum initial state in the magnon mode $b$, the total system population can become almost zero at a desired moment, i.e., $a(\tau),b(\tau)]^{T}=S(t=\tau)[0,b(0)]^{T}\approx(0,0)^{T}$. In practice, the non-Hermitian cavity magnonic system now becomes a promising candidate Ramezani et al. (2014); Longhi (2015); Jin et al. (2016); Ramezani et al. (2016); Huang et al. (2017); Jin and Song (2018); Xu and Jin (2024) for the unidirectional perfect absorber.
Figure 5: Fidelity dynamics about the individual states $|5\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|5\rangle_{b}$ with various $\lambda$, the coupling strength or the dissipation rate of the system, under a $\mathcal{PT}$-symmetric-broken Hamiltonian. Inset: Numerical results for the logarithm of the nonreciprocity $\log_{10}[S_{21}(\tau)/S_{12}(\tau)]$ as a function of $\lambda$. The other parameters are the same as Fig. 3. The initial state of system is $|\psi(0)\rangle=|0\rangle_{a}|5\rangle_{b}$.
In Fig. 5, we demonstrate the fidelity dynamics of the relevant states $|5\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|5\rangle_{b}$ for various $\lambda$. It is found that although the state $|5\rangle_{a}|0\rangle$ is slightly populated, the population on the state $|0\rangle_{a}|5\rangle_{b}$ decreases monotonically with time, with a decay rate roughly proportional to $\lambda$ for $t\leq 0.3\tau$. In particular, when $t=0.1\tau$, we have $\mathcal{F}_{0,5}=0.713$ for $\lambda=0.5$, $\mathcal{F}_{0,5}=0.601$ for $\lambda=1$, and $\mathcal{F}_{0,5}=0.557$ for $\lambda=1.2$. When $t=0.3\tau$, $\mathcal{F}_{0,5}=0.254$ for $\lambda=0.5$, $\mathcal{F}_{0,5}=0.212$ for $\lambda=1$, and $\mathcal{F}_{0,5}=0.199$ for $\lambda=1.2$. When $t\geq 0.8\tau$, the population of the state $|0\rangle_{a}|5\rangle_{b}$ is almost vanishing for various $\lambda$. This result can be verified by the inset of Fig. 5, where the transition coefficients demonstrate a dramatic non-reciprocal relation $S_{21}(\tau)\gg S_{12}(\tau)$ when $\lambda>0.5$, justifying the unidirectional perfect absorbtion in the cavity-magnonic system. Note for $\lambda=0$, $\log_{10}[S_{21}(\tau)/S_{12}(\tau)]=0$, i.e., $S_{21}(\tau)=S_{12}(\tau)$ implies the disappearance of the non-reciprocal behavior, reminiscent of the fact that under the condition of Eq. (35) with $\lambda=0$, the non-Hermitian Hamiltonian in Eq. (37) becomes a Hermitian one Jin and Jing (2025g), i.e., $H(t)=iJ(t)a^{\dagger}b-iJ(t)b^{\dagger}a$. Our results then encompass the bidirectional perfect state transfer in the Hermitian bosonic system as a special case.
V Conclusion
In summary, we propose a systematic and versatile method for manipulating the general bosonic system through a non-Hermitian and time-dependent Hamiltonian. In sharp contrast to conventional protocols, our theory describes the dynamics of non-Hermitian continuous-variable systems in the full Hilbert space through Heisenberg equations, does not ignore the quantum-jump terms, and relies on the gauge freedom of differential manifolds (instantaneous frames) rather than the eigen-spectrum. Consequently, beside realizing unidirectional perfect absorption, the control protocol enables perfect and flexible transfer of arbitrary states between bosonic systems, independent of $\mathcal{PT}$-symmetry of Hamiltonian or coefficient matrix and the presence or absence of exceptional points.
In practice, the system dynamics can be described in the ancillary representations. By rotating from the time-dependent to the stationary ancillary representations, a purely geometric gauge potential emerges to shape the time evolution of the system, which is uniquely determined by the formation of ancillary operators. In the stationary representation, the Hamiltonian’s coefficient matrix can be upper triangularized under Schur’s theorem with the projection operator, which suffices to activate two Heisenberg passages for flexible and perfect state transfers. Along the transitionless passages, arbitrary initial state can be perfectly transferred between the cavity and magnon modes without artificial normalization, including but not limited to Fock states, coherent states, the superposition of Fock states, the cat states, and even thermal states. The resulting passages do not show a clear relation with both $\mathcal{PT}$-symmetry and EPs. Moreover, the cavity-magnonic system can be used as a unidirectional perfect absorber under certain conditions. Essentially, this work extends the framework of our UQC theory, and provides a powerful tool for controlling the non-Hermitian continuous-variable systems.
Acknowledgments
We acknowledge grant support from the National Natural Science Foundation of China (Grant No. U25A20199) and the “Pioneer” and “Leading Goose” R&D of Zhejiang Province (Grant No. 2025C01028).
Appendix A A general recipe for ancillary operators
This appendix provides a brief recipe for constructing the ancillary basis operators $\{\mu_{k}(t)\}$ for a general bosonic systems of $N$ modes as well as the $N\times N$ unitary transformation matrix $\mathcal{M}^{\dagger}(t)$ in Eq. (3). Rooted in the geometric structure of the manifold of $\mu_{k}(t)$’s, the adjoint matrix $\mathcal{M}^{\dagger}(t)$ admits the representation Jin and Jing (2025c, g):
$\mathcal{M}^{\dagger}(t)=\left[\vec{M}_{1}(t),\vec{M}_{2}(t),\cdots,\vec{M}_{N}(t)\right]^{T}$ (42)
with the $N$-dimensional row vectors
$\vec{M}_{1}(t)$ $=(\cos\theta_{1}e^{i\frac{\alpha_{1}}{2}},-\sin\theta_{1}e^{-i\frac{\alpha_{1}}{2}},0,\cdots,0),$ (43)
$\vec{M}_{k}(t)$ $=[\cos\theta_{k}e^{i\frac{\alpha_{k}}{2}}\vec{b}_{k-1}(t),-\sin\theta_{k}e^{-i\frac{\alpha_{k}}{2}},0,\cdots,0],$
$\cdots$
$\vec{M}_{N-1}(t)$ $=[\cos\theta_{N-1}e^{i\frac{\alpha_{N-1}}{2}}\vec{b}_{N-2}(t),$
$-\sin\theta_{N-1}e^{-i\frac{\alpha_{N-1}}{2}}],$
$\vec{M}_{N}(t)$ $=\vec{b}_{N-1}(t),$
where $k$ runs from $2$ to N2N-2. Here $\vec{b}_{k}(t)$ is a $k+1$-dimensional bright vector
$\vec{b}_{k}(t)\equiv[\sin\theta_{k}e^{i\frac{\alpha_{k}}{2}}\vec{b}_{k-1}(t),\cos\theta_{k}e^{-i\frac{\alpha_{k}}{2}}],$ (44)
with $1\leq k\leq N-1$ and $\vec{b}_{0}(t)\equiv 1$. In Eqs. (43) and (44), the time-dependence of the parameters $\theta_{k}(t)$ and $\alpha_{k}(t)$ is treated implicitly for clarity. These parameters may be either time-dependent or time-independent.
With the definition in Eqs. (3) and (42), the rotation from the ancillary operators $\{\mu_{k}(t)\}$ to their stationary version $\mathcal{V}^{\dagger}(t)\mu_{k}(t)\mathcal{V}(t)\rightarrow\mu_{k}(0)$ can be performed by Jin and Jing (2025g)
$\mathcal{V}(t)=V_{\alpha_{1}}V_{\theta_{1}}V_{\alpha_{2}}V_{\theta_{2}}\cdots V_{\alpha_{N-1}}V_{\theta_{N-1}}=\prod_{k=1}^{N-1}V_{\alpha_{k}}V_{\theta_{k}},$ (45)
where
$V_{\alpha_{k}}(t)$ $=e^{-i\frac{\delta\alpha_{k}}{2}\left[b_{k-1}^{\dagger}(0)b_{k-1}(0)-a_{k+1}^{\dagger}a_{k+1}\right]},$ (46)
$V_{\theta_{k}}(t)$ $=e^{-\delta\theta_{k}\left[e^{i\alpha_{k}(0)}a_{k+1}^{\dagger}b_{k-1}(0)-e^{-i\alpha_{k}(0)}b_{k-1}^{\dagger}(0)a_{k+1}\right]}$
with $\delta\alpha_{k}=\alpha_{k}(t)-\alpha_{k}(0)$ and $\delta\theta_{k}=\theta_{k}(t)-\theta_{k}(0)$.
Appendix B Derivation of non-Hermitian Hamiltonian
This appendix provides a detailed derivation of the non-Hermitian Hamiltonian (20) from the Lindblad master equation. We emphasize that the quantum jump terms have been fully retained throughout the derivation Metelmann and Clerk (2015); Wang et al. (2019a) leading to Eq. (20), in contrast to those approaches only regarding the nonunitary evolution part after post-selection Han et al. (2024). In general, the dynamics of an open cavity magnonic system that is coupled to the traveling waves Metelmann and Clerk (2015); Wang et al. (2019a) can be described by
$\frac{d}{dt}\rho=-i[H_{\rm coh},\rho]+\eta\mathcal{L}[c]\rho+\beta\mathcal{L}[a]\rho+\chi\mathcal{L}[b]\rho,$ (47)
where $H_{\rm coh}=\omega_{a}a^{\dagger}a+\omega_{b}b^{\dagger}b+[J\exp(i\varphi)a^{\dagger}b+{\rm H.c.}]$ represents the eigen-energies of the two modes and the coherent coupling between them. The Lindblad superoperators are defined as $\mathcal{L}[o]\rho=o\rho o^{\dagger}-\{o^{\dagger}o,\rho\}/2$, $o=c,a,b$. The first superoperator is about $c\equiv ua+\exp{(i\Theta)}vb$ with weight factors $u$ and $v$. It implies the cooperative interactions between the two modes and the traveling wave with a damping rate $\eta$. $\Theta=0$ or $\Theta=\pi$ is determined by the propagation direction of the traveling waves. The second and the third superoperators are associated with the individual dissipations of the modes $a$ and $b$, with the damping rates $\beta$ and $\chi$, respectively.
Under Eq. (47), the Schrödinger-picture operator $\mathcal{O}_{S}$ is connected to the Heisenberg-picture operator $\mathcal{O}_{H}(t)$ as
${\rm Tr}\left[\mathcal{O}_{S}\dot{\rho}(t)\right]$ (48)
$={\rm Tr}\left[\mathcal{O}_{S}\left(-i[H_{\rm coh},\rho]+\eta\mathcal{L}[c]\rho+\beta\mathcal{L}[a]\rho+\chi\mathcal{L}[b]\rho\right)\right]$
$={\rm Tr}\left[\left(i[H_{\rm coh},\mathcal{O}_{S}]+\eta\mathcal{L}^{\dagger}[c]\rho+\beta\mathcal{L}^{\dagger}[a]\rho+\chi\mathcal{L}^{\dagger}[b]\right)\rho\right]$
$={\rm Tr}\left[\dot{\mathcal{O}}_{H}(t)\rho(0)\right],$
where the Hermitian conjugate superoperator $\mathcal{L}^{\dagger}[o]\equiv o^{\dagger}\rho o-\{o^{\dagger}o,\rho\}/2$. The derivation from the second line to the third one in Eq. (48) has used the cyclic property of the density trace. Accordingly, the time derivative of $\mathcal{O}_{H}(t)$ can be expressed as
$\frac{d}{dt}\mathcal{O}_{H}(t)=i[H_{\rm coh},\mathcal{O}_{S}]+\eta\mathcal{L}^{\dagger}[c]\rho+\beta\mathcal{L}^{\dagger}[a]\rho+\chi\mathcal{L}^{\dagger}[b].$ (49)
Using Eq. (49), the dynamics of the modes $a$ and $b$ can be obtained as
$\frac{d}{dt}a(t)$ $=-i\omega_{a}a-\gamma_{a}a-\left(iJ+e^{i\Theta}\Gamma\right)b,$ (50)
$\frac{d}{dt}b(t)$ $=-i\omega_{b}b-\gamma_{b}b-\left(iJ+e^{i\Theta}\Gamma\right)a,$
under the conditions of $\Gamma=\eta uv$, $\gamma_{a}=\eta v^{2}+\beta$, and $\gamma_{b}=\eta u^{2}+\chi$. The system dynamics described by Eq. (50) is equivalent to that governed by the Hamiltonian (20) with $\varphi_{a}=0$.